A Bibliography of Publications by, and about, Walter Gautschi

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 August 2025 Version 0.17

Title word cross-reference

$24.50 [G.70]. $49 [I.91]. $69.95 [Ano03]. cos(x)/x [Gau66a, GK70b]. e^{x}/x [Gau66a, GK70b]. f(z)/z [Gau66c]. ∫_{1}^{∞} e^{−xt}t^{−n} dt [Gau59a]. n [Gau55, Gau59a]. S [GM97]. sin(x)/x [Gau66a, GK70b]. W [Gau11b].

-functions [Gau11b]. -orthogonality [GM97]. -ter [Gau55]. -th [Gau55].

16 [JW08]. 16-20.07.2007 [JW08]. 1734 [Gau08e]. 1943-1993 [Len97]. 1975 [Ask75]. 1983 [GW84]. 1993 [Zah94b].

20.07.2007 [JW08]. 221 [Gau64c]. 222 [Gau64a, Gau64d, Gau64e]. 224 [Ano89]. 236 [Gau64b, Gau65b]. 25 [Ano89]. 259 [Gau65a, Jan77]. 282 [Gau66a]. 292 [Gau66b, Gau69e].

331 [Gau68a]. 363 [Gau69a, Gau69b, Köl72].

420 [Ell74]. 471 [Gau73a].

50th [Gau94b]. 521 [Gau77a]. 542 [Gau79a].

65th [Ric94, Won90, Zah94a]. 6th [JW08].

7 [Wei95]. 726 [Gau94a, Gau98b]. 793 [Gau99a].

abstract [Gau62b]. accident [Gau18d]. ACM [Gau64b, Gau64a, Gau69a, Gau73a]. Adaptive [GG98, GG00]. advanced [Ask75]. Advances [Gau76]. aerodynamics [Gau12b, Gau13a]. after [Gau81c, Gau12a]. agli [Gau56b]. Airy [Gau02b]. Alexander [Gau83d, SDGT81]. Algebra [Hig94, Gau02c]. algebraic [Gau73c, GN88, Gau12b, Gau13a]. algebraic/logarithmic [Gau12b, Gau13a]. Algorithm [Gau64b, Gau64a, Gau69a, Gau73a, Gau16a, KD71, FG86, PPS18, Ell74, Gau64c, Gau64d, Gau64e, Gau65a, Gau65b, Gau66a, Gau66b, Gau68a, Gau69b, Gau69e, GK70b, Gau77a, Gau79a, Gau94a, Gau98b, Gau99a, Jan77, Köl72, KD71, Sko75]. Algorithms [Gau11e]. always [Gau93a]. Anal [Wei95]. Analyse [Gau53a]. Analysis [Gau81a, Gol84, Won90, Gau72, Gau72b, Gau97b, Gau97c, Gau02c, Gau10c, GM24, Ste99, SB93, Gau53a, Ano03]. Analytic [Gau59b, Gau17b, GV83, GL90]. Angelitch [G.70]. Anniversary [Gau94b]. Anomalous [Gau77b]. Application [GM85a, GM85b, Ask75, Gau69c, Gau12b, Gau13a, Qi09]. Applications [GGO99, Gau85a, Gau85b, Gau96b]. Applied [Ano89, Ano03, Ano98, Gri05b, JW08]. approach [KR15]. appropriate [Gau54a]. Approximation [Ell96, Gau04b, Gau06c, GMR11, FGM87, Gau18a, Zah94b, Seg06]. approximations [Gau84a, GM86, Gau02a]. April [Ask75]. Architecture [Roj94]. Argument [Gau62a]. Arguments [Gau65a, Jan77]. Arising [Gau69d]. Askey [Gau18b]. Aspects [Gau67a, Gau72a, Gau73b]. Asymptotic [Gau53b, Gau53c, Won90]. asymptotics [GG08]. attempt [Gau08e]. Attenuation [Gau72, Gau72b]. August [Gau94b]. Aumann [G.70].

B [BF81, Gau17c]. B-spline [Gau17c]. Bailey [Gau96a]. Bartels [Ano03]. Barycentric [Gau01a]. based [Gau61a]. Bauer [G.70]. Bedingung [Gau56a]. before [Gau12a]. behavior [Gau83c]. Behaviour [Gau53b, Gau53c]. beim [Gau54a]. Bemerkung [Gau56a]. Bemerkungen [För82]. Bemessung [Gau54a]. Berlin [G.70]. Bernoulli [Gau08e]. Bernstein [Ano89, GN89, Gau10a]. Berrut [Gau01a]. Bessel [Sko75, Gau64b, Gau65b, GS78, Gau02b, Kah6x]. best [Qi09]. Beta [Gau64a, Gau64d, Gau64e]. between [Gau02c]. bibliography [Gau07]. Binet [GM18]. Binet-type [GM18]. Birkhauser [I.91]. Birthday [Won90, Gau83d, Ric94, Zah94a]. Bivariate [Gau59b]. Bjørstad [Gau96a]. Blue [Gau79d]. Book [Ano03, Ell96, G.68, G.70, Gau96a, Gau00a, I.91, Len97, Ste99]. Bose [Gau93b]. Boston [I.91]. Bounds [GI88, GV83, Gri05b, Kah6x, Qi09]. British [Gau94b]. Bulirsch [Ano03, G.70].

C [Ano03]. Calculation [Gau99b, Gau56a, Gau70c]. cardinal [Gau01a, Gau17c]. career [Gau18d]. case [Gau11b]. Center [Ask75]. Century [Lid01, Gau94b, Len97]. Certain [Gau61b, GK70a, Gau91b]. Certification [Gau64e, Gau65b, Köl72]. challenging [Gau08c]. Chebyshev [AG75, För82, GY74, Gau75b, Gau76, GM77, GTV90]. Chebyshev-type [GM77, GY74, Gau75b, AG75, För82]. check [Gau83b]. chemistry [Gau84b]. Christoffel [BF81, Gau68b, Gau81a, Gau81c, Gau81d, Gau21]. Chronicle [Lid01]. circle [CGR90, Gau06a]. class [Gau54b, Gau91a, GN96, QG08, Qi09]. classes [Gau10b]. classical [Gau02c]. cm [G.70, I.91]. coefficients [Gau54b, Gau70c, Gau09d]. Coerror [Gau77a, Gau77c, Gau16a]. Collaboration [Gau11c]. Columbia [Gau94b]. Commentaries [BS13, BS14a, BS14b]. commentary [Gau08e]. completely [QG08, Qi09]. Complex [Gau69a, Gau70a, Gau05d, PPS18, Wei95, Gau69b, Köl72]. Componentwise [Hig94]. Computation [Ell96, Gau61b, Gau61c, Gau66c, Gau70a, Gau02b, Gau04b, Gau05d, GMR11, Len97, Seg06, Wei95, Gau63b, GK70a, GS78, Gau93b, Gau94b, Gau96b, GGO99, Gau11b, GM24, Zah94b]. Computational [Ano89, Gau67a, Gau75a, Gau79b, Gau00a, Gau06c, Won90, Gau94b, Len97]. compute [Gau08e]. Computer [Roj94]. Computing [CGM86a, CGM86b, GW87, GZ95, GW01, Gau05a, Gau06b, Gau96a]. Condition [Gau69d, Gau72c, Gau79c, Gau83a, GI88, Gau56a, Gau56b, Gau73c, Gau84c]. Conditioned [Gau75d, Gau11d]. conditioning [CGR90]. conditions [Gau78b]. condizione [Gau56b]. Conference [Won90, GGO99, Zah94b]. confluent [Gau62c, Gau63a, Gau78a, Gau02a]. congress [JW08]. conjecture [Gau18b, JB09]. Conjectured [Gau87, GL07, Gau08d, Gau09c, Gau09b, Gau11e, Kou07]. conjectures [GH19, LR14]. constant [Gau54b]. Construction [Gau68b, Gau70b, GM97]. Constructive [Gau85a]. contact [Gau91b]. contemporary [Ano98]. continued [Gau77b, Gau83c]. Contour [GTV90]. control [KR15]. convergence [Gau77b, Gau83c, Gau92, GL96]. convergent [Gau91a, Gau91b, Gau05b]. Correction [Gau19]. coth [GKM87]. coth-kernel [GKM87]. Coulomb [Gau66b, Gau69c, Gau69e]. coupled [XDY16].

D [Gau96a]. D1 [Gau68a]. Daniel [Gau08e]. David [Gau96a]. Debate [Roj94]. decaying [Gau05a]. December [Zah94b]. definite [FG86]. dell’errore [Gau75e]. densely [Gau05a]. Derivatives [Gau66c, GK70b, GK70a, GM24, Gau66a]. diagonal [FG86]. dicembre [Ano98]. Difference [Gau62b]. Differential [G.68, AG62, Gau55, Gau61a, Gau75e, Gri05b, Gri05a, HL99]. Differentialgleichungen [Gau55]. differenziali [Gau75e]. Diffusion [Gau62a]. dimensioning [Gau54a]. Dirac [Gau93b]. Discrete [Gau84a, Gau18a]. discretization [XDY16]. discretizations [BHvdV06]. distributions [Gau84a, GM86]. divisors [Gau13c]. Donald [G.68]. doppi [Gau56b]. double [Gau56b, GL91]. down [Gau18a]. during [Gau81c].

E. [BF81]. edge [BHvdV06]. edited [Ell96, Gau96a]. Edition [Ano03]. Editor [Len97, G.68]. Efficient [Gau70a]. Einstein [GM85a, GM85b, Gau93b]. element [BHvdV06]. Elementary [Gau59c]. elements [Gau83c]. empirical [Gau17b]. end [GL90, GL91]. energy [XDY16]. equation [KR15]. Equations [Gau62b, AG62, BHvdV06, Gau55, Gau61a, Gau73c, Gau75e, Gri05b, Gri05a, HL99, XDY16, G.68]. equazioni [Gau75e]. Erratum [Ano89, Gau13a, Gau17a, Wei95]. Error [Gau61c, Gau64f, Gau69a, Gau69b, Gau70a, GV83, XDY16, Gau55, GK70a, Gau75e, Gri05b, Köl72, Wei95, Gau53d]. estensione [Gau56b]. Estimates [Gau75c, Gau53d, XDY16, Gau75e]. estremo [Gau56b]. Euler [Fel07, Gau08e, Gau08a, Gau08b]. Evaluation [Gau77c, Gau16a, GKM87, Gau08c]. Excellence [Lid01]. exercise [Gau05b]. Exercises [Tow16, Gau16b]. Expansions [GHT03]. Experimental [Gau11a]. Explicit [SW19]. Exponential [Gau59a, GC64, Gau73a, GHT03, Gau15, Gau22, Gri05b]. exponentially [Gau05a]. extended [Gau92, GL96]. extension [Gau56b]. extensions [Ker83, QG08]. extreme [Gau56b].

F [G.70]. factors [Gau72, Gau72b]. Family [GR88]. February [Gau08e]. Fehler [Gau55]. Fehlerabschätzungen [Gau53d]. Fermi [GM85a, GM85b, Gau93b]. Festschrift [Ell96, Zah94b]. field [Pom94]. finite [BHvdV06, FGM87]. First [Gau64b, Sko75, Gau65b]. Foreword [Ric94]. Form [Gau79c, FG86]. Formulae [CGM86a, CGM86b, Gau81d, GR88, Gau04a, Ano89, Gau83b, GN88, GN89, Gau91c, GN96, GM97, Gau00b, Gau00c, Gau01a, Gau02b, Gau09a, JB09]. Formulas [AS64, Gau68a, Gau68b, AG75]. Fourier [Gau70c, Gau72, Gau72b]. fraction [Gau77b]. fractions [Gau83c]. Frank [Won90]. frequency [SW19]. Fresnel [Gau64f]. Freud [Gau17b]. Function [Gau59c, Gau61c, Gau64a, Gau64f, Gau69a, Gau69b, Gau70a, Gau74a, Gau74b, Gau77a, Gau77c, Gau05d, Gau16a, Alz03, Gau64e, GS78, Gau82b, GW87, Gau05b, GLP98, Ker83, Köl72, Wei95]. Functions [AS64, Gau59b, Gau62a, Gau64b, GC64, Gau65a, Gau66b, Gau75a, Gau79a, Gau79b, GM85a, GM85b, Gau99a, Gau99b, Jan77, Sko75, Ano89, Ask75, Gau63b, Gau64c, Gau64d, Gau65b, Gau69c, Gau69e, Gau77b, GV83, GN88, GN89, GL90, GN96, Gau98a, Gau00b, GGL00, GW01, Gau01b, Gau02b, Gau02a, GHT03, Gau05a, GG08, Gau09a, Gau10b, Gau10c, Gau11b, Gau14, Gau17c, Gau21, Gau22, GM22, Kah6x, QG08, Qi09, BL01].

G [G.70]. Gamma [Gau59c, Gau74a, Gau74b, Gau79a, Gau79b, Gau99b, Alz03, Gau82b, Gau98a, GHT03, GLP98, Ker83, Gau64c]. Gatteschi [GG08]. Gauss [Ano89, CGM86a, CGM86b, Gau68b, Gau81d, GN88, GR88, GN89, GTV90, GL90, GL91, Gau94a, GM97, Gau98b, Gau99a, Gau00b, Gau00c, Gau02a, Gau04a, Gau05c, Gau06a, Gau09a, Gau10b, Gau14, Gau16a, JB09, PPS18]. Gauss-Type [Gau94a, Gau98b, Gau06a]. Gaussian [Gau68a, Gau70b, Gau70c, GV83, Gau83b, GM85a, GM85b, Gau91a, Gau02b, Gau21]. Gautschi [Ano03, I.91, Len97, Zah94b, Alz03, BHvdV06, BS13, BS14a, BS14b, Dav04, GLP98, Gri05b, Gri05a, HL99, JB09, Ker83, Kou07, KR15, LR14, QG08, Qi09, Ric94, Seg06, SW19, Ste99, Tow16, XDY16, Zah94a, Zah94c, Ell96]. Gautschi-type [Gri05b, Gri05a, HL99, SW19, XDY16]. Gene [Gau02c]. Generalized [Gau04a, Gau93b, GW01, Gau09a, GM21, JB09]. Generating [Gau82a, Gau94a, Gau98b, Kah6x]. German [I.91, För82, Gau53a, Gau53d, Gau54b, Gau54a, Gau55, Gau56a, Gau72a]. Germany [GGO99]. gewöhnlicher [Gau55]. Gilbert [Gau96a]. global [Gau75e]. globale [Gau75e]. Golub [Gau02c]. Gordon [XDY16]. GQRAT [Gau99a]. Gradimir [GMR11, Gau11c]. Grammel [Gau53d, Gau53d]. graphical [Gau53a, Gau53d, Gau54a]. graphischen [Gau53d, Gau54a]. graphischer [Gau53a]. Graphs [AS64]. Greenspan [G.68]. guided [Gau07].

H [G.70, Gau96a, Gau02c, KD71]. Half [Gau16a, Gau91c, Gau94b, Len97]. half-century [Gau94b, Len97]. half-infinite [Gau91c]. Half-Range [Gau16a]. Handbook [AS64, BL01]. Hardy [Gau05b]. Harmonic [Gau74a, Alz03]. Heinz [I.91]. Hermite [Gau87, GW01, Gau14, Gau16a]. Hidden [Ell74]. Hidden-Line [Ell74]. High [Gau00c, Gau09a, Gau14, Gau11b]. High-order [Gau00c, Gau09a]. High-precision [Gau14, Gau11b]. Hilbert [GW87, GW01]. Hilfsmittel [G.70]. historical [Gau05c]. Honor [GMR11, Won90, Ell96, Zah94b]. Honoring [SDGT81]. Horst [Gau96a]. hypergeometric [Gau02a].

ICIAM [JW08]. ideas [Ano98]. II [Gau53c, Gau63a]. III [G.70, Gau78a]. inaccuracies [Gau54a]. Incomplete [Gau59c, Gau64a, Gau79a, Gau79b, Gau99b, Gau98a, GHT03, Gau64d, Gau64e]. industrial [JW08]. Inequalities [Gau59c, Gau74b, LR14, GL07, Gau09c, Gau09b, Gau11e, GLP98, Ker83, QG08]. Inequality [Gau74a, Gau87, Alz03, Gau08d, Gau10a, Kou07, Qi09]. Infinite [Gau75b, Gau91c]. influence [BF81]. Ingenieurs [G.70]. initial [SW19]. Instability [Gau62b]. Institute [GGO99]. Integral [GC64, GTV90, Gau06b, Gau16a, Gau59a, Gau79d, GHT03, Gau08c, Gau22, GM24]. integrali [Gau56b]. Integrals [Gau61b, Gau61c, Gau64f, Gau73a, Gau77a, Gau77c, Gau56b, GKM87, Gau93b, Gau05a, Gau11b, Gau15]. Integration [Gau55, Gau53a, Gau53d, Gau54a, Gau61a, Gau95, Gau12b, Gau13a, Gau00a]. Integrationsmethoden [Gau53a]. Integrationsverfahren [Gau53d, Gau54a]. integrator [Gri05b, XDY16]. integrators [SW19]. interest [Gau84b]. International [JW08]. interplay [Gau02c]. interpolants [Gau01a]. Interpolation [Gau59b, Gau87, Gau12a, Gau92, GL96, Gau08e]. Intervals [Gau75b, FGM87, Gau91c]. Interview [Dav04]. Introduction [SB93, Zah94a, Gau97c, Ste99, Ano03]. Invented [Roj94]. Inverses [Gau75c, Gau06b, Gau62c, Gau63a, Gau78a]. Inversion [Gau69d]. Involving [Gau83a, GM85a, GM85b, Gau11a]. Ismail [Gau18b]. Italian [Gau56b, Gau75e].

J [Ano03, G.70, Gau96a, Wei95, Won90]. Jacobi [Gau11e, Gau17a, Gau19, GW87, Gau87, GN88, Gau00b, GL07, Gau08d, Gau09c, Gau09a, Gau10a, Gau09b, Gau12c, Gau18c, Kou07, LR14]. James [Gau79d]. Jean [Gau01a]. Jean-Paul [Gau01a]. John [G.68, Gau96a]. Journal [Ano89]. June [GW84].

kernel [GKM87]. Kershaw [GLP98, Qi09]. Kind [Gau64b, Sko75, Gau65b]. Klasse [Gau54b]. Klein [XDY16]. Koeffizienten [Gau54b]. konstanten [Gau54b]. Kontorovich [Gau06b]. Kronrod [Ano89, CGM86a, CGM86b, GN88, GR88, GN89, Gau05c]. Kummer [Gau77b]. Kunzi [G.70]. Kutta [Gau55, Gau55]. Kutta-Verfahrens [Gau55].

L [G.70, Gau79d, GW84]. Lagrange [Gau92, GL96, Gau12a]. Laguerre [Gau00b, GW01, Gau09a, Gau14]. Lambert [Gau11b]. Lanczos [PPS18]. Laplace [Gau69d]. large [Gau59a]. Larger [Gau65a, Jan77]. largest [GL07, Gau08d]. Layne [Gau96a]. Lebedev [Gau06b]. Lecture [Gau08a]. Lectures [RG90, I.91]. Legendre [Gau79d, Gau65a, Jan77]. length [Gau54a]. Leonhard [Fel07, Gau08a, Gau08b]. Leopardi [Kou07]. Letessier [Gau18b]. letter [Gau08e]. Life [Gau08b, Gau08a]. Lifetime [Gau81c]. like [Gau83a, Gau11d]. Line [Ell74]. Linear [Gau59b, Gau62b, Hig94, KD71, Gau54b, Gau02c, Gau13c]. linearen [Gau54b]. Littlewood [Gau05b]. Lobatto [GL90, GL91, Gau00c, Gau04a, Gau09a, JB09]. logarithmic [Gau10b, Gau12b, Gau13a]. logarithmically [QG08, Qi09]. logarithms [Gau08e]. look [Gau22]. Lower [GI88]. Ludwig [Gau53d, Gau54a]. Luigi [GG08]. Luke [GW84].

M [Ell96, Gau83d, SDGT81]. Macdonald [Gau05d]. Madison [Ask75].

Man [Gau08b, Gau08a]. March [Ask75, GGO99]. Marchenko [GM21]. Markov [Gau18a]. Mascagni [Gau96a]. Mathematical [AS64, BL01, GGO99]. Mathematics [Ano89, Ano03, Ask75, Gau11a, Len97, Ano98, BF81, Gau94b, JW08, RG90, Gau94b, I.91, Len97]. Mathematische [G.70]. MATLAB [Tow16, Gau16b, Gau05e, Gau06c]. Matrices [Gau53b, Gau53c, Gau75c, Gau75d, Gau83a, GI88, CGR90, FG86, Gau62c, Gau63a, Gau78a, Gau11d]. Matrix [Gau69d]. Maxwell [BHvdV06]. May [GW84]. Mean [Gau74a, Gau74b, Alz03, Gau92]. measure [GM21]. Measurements [Lid01]. Meissner [Gau53d, Gau54a]. Meissner-Ludwig [Gau53d, Gau54a]. memoriam [GW84]. method [Gau54a, Gau55, Gri05a, HL99]. Methods [Gau75a, Gau06c, AG62, Gau53a, Gau75e]. metodi [Gau75e]. Michael [Gau96a]. Milovanović [GMR11, Gau11c]. Minimal [Gau81b, Gau69c]. Minit [KD71]. modifications [Gau13c]. Modified [Gau70b, GS78]. Moment [FGM87]. Moment-preserving [FGM87]. Moments [Gau70b, Gau97a, Gau86]. monotonic [QG08, Qi09]. Monotonicity [Gau17b, Gau18b]. multi [SW19]. multi-frequency [SW19]. multiple [GL90]. My [Gau11c, Gau07, Gau18d].

NBS [Lid01]. NBS/NIST [Lid01]. nearby [Gau13b]. nearly [FG86]. necessaria [Gau56b]. necessary [Gau56a, Gau56b]. nei [Gau75e]. Neutralizing [Gau13b]. ninetieth [Gau83d]. NIST [Lid01]. no [Wei95]. Nonexistence [Gau75b]. Nonlinear [G.68, SW19, XDY16]. nonstandard [Gau09d]. Norm [Gau75c]. Note [Gau59b, GTV90, Gau99b, Kah6x, Gau05c, Gri05a, Gau01a]. notwendigen [Gau56a]. novembre [Ano98]. Number [GI88, Pom94]. Numer [Gau11e, Wei95]. Numerical [AG62, Gau61a, Gau67b, Gau69d, Gau73b, Gau81a, Gau97b, Gau97c, Gau05d, Gau12b, Gol84, Hig94, I.91, Ste99, Ano03, Gau55, Gau78b, Gau84c, GKM87, Gau95, Gau01b, Gau02c, Gau08c, Gau10c, Gau13b, RG90, SB93, G.68, Gau72a, Gau13a]. Numerik [Gau72a]. numerische [Gau55].

Oberwolfach [GGO99]. occurring [Gau91b]. Olver [Won90]. One [Gau65a, Jan77, Gau56b, Gau75e]. one-step [Gau75e]. only [Gau62b]. Optimal [AG75, För82, GM77, KR15]. optimalen [För82]. Optimally [Gau75d, Gau11d]. Order [Gau62b, Gau55, Gau00c, Gau09a, Gri05b, Gri05a, HL99, SW19]. Orders [Gau05d]. ordinarie [Gau75e]. ordinary [AG62, Gau55, Gau61a, Gau75e]. Ordnung [Gau55]. Orthogonal [Gau72c, Gau82a, Gau83a, Gau85a, Gau94a, Gau96b, Gau99c, Gau04b, Gau05e, Gau06c, Gau11a, Gau16b, GM21, GM22, FG86, Gau81b, Gau82b, Gau84b, Gau85b, Gau86, Gau89, Gau93a, GZ95, Gau98b, GGO99, Gau05a, Gau09d, Gau13c, Gau15, Gau17c, Gau18e, GH19, Gau22, Seg06, Tow16]. orthogonality [GM97]. ORTHPOL [Gau94a, Gau98b]. oscillating [Gau05a]. oscillatory [Gri05b, Gri05a, HL99, SW19]. Ostrowski [Gau83d, SDGT81].

P [G.70]. Package [Gau94a, Gau98b]. pages [Ano03]. paper [Gau79d]. Parallel [Gau96a]. Pastur [GM21]. Paul [Gau01a]. Perturbation [Hig94]. perturbations [Gau86]. Petter [Gau96a]. Philip [Gau95]. physical [BF81]. Picone [Gau56a, Gau56a, Gau56b]. plate [Gau91b]. Plotting [Ell74]. points [GL90, GL91]. polynomial [Gau79d]. Polynomials [Gau72c, Gau79c, Gau82a, Gau82b, Gau83a, Gau87, Gau94a, Gau04b, Gau06c, Gau11a, Gau15, Gau17c, Seg06, Tow16, Gau61a, Gau78b, Gau81b, Gau84c, Gau84b, Gau85a, Gau85b, Gau86, Gau89, Gau93a, GZ95, Gau96b, GN96, Gau98b, GGO99, Gau99c, Gau05a, Gau05e, GL07, Gau08d, Gau09c, Gau10a, Gau09b, Gau09d, Gau11e, Gau12c, Gau13c, Gau16b, Gau17b, GM18, Gau18b, Gau18e, Gau18c, GH19, GM21, Gau22, GM22, Kou07, LR14, Gau17a, Gau19]. positive [FG86]. Power [Gau79c]. Powers [Gau53b, Gau53c]. pp [G.70, I.91]. pp. [G.68]. practical [Gau72, Gau72b]. preceding [Gau79d]. precision [Gau09d, Gau11b, Gau14]. Preface [GMR01]. Presence [Gau67b, Gau12b, Gau13a]. preserving [FGM87]. Price [G.70, I.91]. probability [GM21]. problem [Gau18a]. problems [Gau91b, Gau97a, SW19]. Procedure [Gau79b]. procedures [Gau53d]. proceedings [Ask75, Zah94b]. processing [Gau96a]. Program [Ell74]. Programming [KD71]. Progress [Gau18d]. propagation [GK70a]. properties [CGR90, Gau17b]. Prudnikov [GM22]. Publications [Lid01, Zah94c]. Purdue [Zah94b].

Quadratur [För82]. Quadrature [CGM86a, CGM86b, Gau67b, Gau68a, Gau68b, Gau70b, Gau81c, Gau81d, GM85a, GM85b, GR88, GTV90, Gau91c, Gau94a, Gau99a, GGL00, Gau05d, Gau06c, Gau16a, AG75, Ano89, GG98, GG00, Gau70c, Gau76, GV83, Gau83b, GN88, GN89, GL90, Gau91a, GL96, GN96, Gau97a, GM97, Gau98b, Gau99c, Gau01b, Gau02b, Gau02a, Gau05c, Gau06a, Gau10b, Gau13b, Gau14, PPS18]. Quadratures [GY74, Gau75b, För82, GM77, GL91, Gau21]. Questions [Gau78b, Gau84c].

R [Ano03, G.70, Gau96a]. R. [Ell96, G.70]. Rabinowitz [Gau95]. Radau [GL90, GL91, Gau00b, Gau04a, Gau09a, JB09]. Ramanujan [GM24]. Range [Gau16a, Gau12c, Gau17b, Gau17a, Gau19]. Rational [Gau99a, GGL00, Gau01b]. Ratios [Gau64a, Gau64d, Gau64e, Gau77b, GS78]. real [Gau83c]. reciprocal [Gau82b]. Recognition [Gau81c]. Recurrence [Gau67a, Gau72a, Gau73b, Gau81b, Gau93a, Gau09d, Kah6x]. recurrences [Gau69c]. Recursive [Gau61b, Gau61c, Gau63b, GK70a, Gau99b]. Refinements [QG08]. reflections [Gau18d]. Regular [Gau66b, Gau69e]. rekurrenter [Gau72a]. Related [GC64, Gau78b, Gau84c, GN96, Gau02b, Gau05a, Gau06a]. Relating [Gau59c]. relation [Gau93a]. Relationen [Gau72a]. Relations [Gau67a, Gau72a, Gau73b, Gau81b]. relative [GM21, Gau22, GM22]. Remainder [GTV90, GL90]. Remark [Ell74, Gau69e, GK70b, Gau98b, Gau11e, Jan77, KD71, Sko75, Gau56a]. Remarks [För82]. Repeated [Gau61c, Gau77a, Gau77c, Gau13c, Gau16a]. repository [Gau18e, Gau21]. Representation [GTV90]. Research [Ask75, GGO99]. respect [Gau82b, Gau05a, Gau15, Gau17c]. results [Gau17b]. Review [Ano03, G.68, G.70, I.91, Seg06, Gau00a, Tow16]. Reviews [Ell96, Gau96a, Len97, Ste99]. Revisited [GG98, GG00]. Robert [Gau96a]. Roma [Ano98]. Routines [Gau94a, Gau98b, Gau10b]. Rule [GTV90]. Rules [Gau70b, Gau94a, Gau70c, GL90, Gau98b, GGL00, Gau06a, Gau14]. Runge [Gau55, Gau55]. Rutishauser [G.70, I.91].

S [Gau96a]. S13 [Gau73a]. S14 [Gau79a]. S15 [Gau69b, Gau77a]. S16 [Gau65a]. S17 [Sko75, Gau64b, Gau65b]. S22 [Gau69e]. Samelson [G.70]. Sauer [G.70]. scaled [Gau11d]. scheme [BHvdV06]. Schreiber [Gau96a]. Schrittlänge [Gau54a]. sciences [BF81]. scientific [Gau96a]. Second [Gau62b, Gri05b, Gri05a, HL99, Qi09, SW19]. Second-Order [Gau62b, Gri05b, Gri05a, HL99, SW19]. Selected [BS13, Lid01, BS14a, BS14b]. semicircle [Gau89]. seminar [Ask75]. sensitivity [Gau86]. Series [GM85a, GM85b, Gau91a, Gau91b, Gau05b]. several [FG86]. sharp [Gau10a]. SIAM [Wei95]. sieve [Pom94]. Simon [Gau96a]. simultaneous [FG86]. SINC [Gau01a]. SINC- [Gau01a]. since [Gau98a]. singular [GKM87]. singularities [Gau12b, Gau13a, Gau13b]. Singularity [Gau67b]. slowly [Gau91a, Gau91b, Gau05b]. Small [Gau62a]. Sobolev [GZ95]. Software [Gau06c, Gau18e, Gau21]. Solutions [G.68, Gau69c, Gau81b, Gau16b, Tow16]. Some [Gau59c, Gau74b, Gau85b, Ker83, Gau84b, Gau11b, Gau18d]. Sons [G.68]. space [XDY16]. spaces [GZ95]. Special [Gau75a, Gau08a, Ask75, Gau63b, GG08, Gau10c]. spectral [CGR90, XDY16]. spherically [Gau84a, GM86]. spiral [Gau10c]. Spline [GM86, FGM87, Gau17c]. sponsored [Ask75]. Springer [Ano03, G.70]. Springer-Verlag [G.70]. square [Gau12b, Gau13a]. stable [Gau90, Gau93a]. Standards [Lid01]. Stenger [GH19]. step [Gau54a, Gau75e]. stepping [BHvdV06, KR15]. Stieltjes [GN96]. Stime [Gau75e]. Stoer [Ano03, G.70]. Studies [Gol84]. study [GK70a, GN88, Gau11b]. Sturm [LR14]. Sub [Gau12c, Gau17b, Gau17a, Gau19]. Sub-range [Gau12c, Gau17b, Gau17a, Gau19]. subrange [Gau18c]. Successive [Gau66c]. Summation [GM85a, GM85b, Gau91a]. Supplement [CGM86b, GM85b]. Survey [Gau75a, Gau81d, Hig94]. Switzerland [JW08]. symmetric [FG86, Gau84a, GM86]. Symposium [Gau94b]. Systemen [Gau54b]. systems [Gau54b, Gau90]. Szabo [G.70]. Szego [Ano89, GN89].

T [G.70, Gau96a]. Tables [AS64, Gau70c]. Technology [Lid01]. ter [Gau55]. Term [Gau67a, GTV90, Gau69c, Gau81b, GL90]. Texts [Ano03]. th [Gau55]. their [Gau91a, Gau06b, GL07, GG08, Gau11b, GM18]. Theodorus [Gau10c]. theorem [Gau06a, LR14]. theorems [Gau06a]. theoretical [Gau84b]. Theory [Hig94, Gau85a, Gau18a, GH19, Ask75]. Third [Ano03]. Three [Gau67a, Gau69c, Gau81b]. Three-Term [Gau67a, Gau69c, Gau81b]. time [BHvdV06, KR15]. time-stepping [KR15]. top [Gau18a]. top-down [Gau18a]. Torczon [Gau96a]. Torino [Ano98]. tour [Gau07]. Transform [Gau69d, GW87, GW01]. transformation [FG86]. Transforms [Gau06b]. Translated [Ano03, I.91]. treatment [GLP98]. tribute [Gau02c]. Tricomi [Ano98, Gau98a]. trigonometric [Gau61a]. Tschebyscheff [För82]. Tschebyscheff-Typ [För82]. Turán [GM97, Gau14]. Turán-type [GM97]. two [Gau10b]. Typ [För82]. Type [GY74, Gau75b, Gau94a, AG75, Ano89, För82, GN89, GM97, Gau98b, Gau06a, GM18, GM22, GLP98, Gri05b, Gri05a, HL99, SW19, XDY16, GM77].

ultraspherical [Gau18b]. Ungenauigkeiten [Gau54a]. Unified [GLP98]. University [Ask75]. use [Gau01b].

V [Ell96, Gau96a, GMR11, Gau11c]. Value [Gau74b, SW19]. values [Gau59a]. Vancouver [Gau94b]. Vandermonde [CGR90, Gau62c, Gau63a, Gau75c, Gau75d, Gau78a, Gau83a, GI88, Gau90, Gau11d]. Vandermonde-like [Gau83a, Gau11d]. Variable [Gau09d]. Variable-precision [Gau09d]. variation [Gau56a]. Variationsrechnung [Gau56a]. Verfahrens [Gau55]. Verlag [G.70]. via [LR14]. Viewpoint [Roj94]. Virginia [Gau96a]. Volume [BS13, BS14a, BS14b].

W [Ano03, Won90]. W. [Ker83]. Walter [Ell96, I.91, Seg06, Zah94b, BS13, BS14a, BS14b, Dav04, Gau11e, Len97, Ric94, Ste99, Zah94a, Zah94c]. Watson [Gau96a]. Wave [Gau66b, Gau69c, Gau69e, KR15]. weight [Ano89, GW87, GN88, GN89, GN96, Gau00b, GW01, Gau05a, Gau09a, Gau10b, Gau14, Gau17c, Gau22, GM22]. weighted [AG75]. Wendel [QG08]. Who [Roj94]. Wiley [G.68]. Wisconsin [Ask75]. Wisconsin-Madison [Ask75]. Witzgall [Ano03]. Work [Gau81c, BF81, Gau95, GG08]. Works [BS13, Gau08b, BS14a, BS14b, Gau08a].

x [G.68]. xv [I.91]. xviv [G.70].

York [Ano03, G.68]. Yudell [GW84].

Zahar [Ell96]. zeichnerischen [Gau54a]. zero [Gau08d]. Zeros [Gau87, Gau89, GL07, GG08, Gau09c, Gau09b, Gau11e, Gau17b, GM18, Gau18b, Gau18c, LR14]. zur [För82, Gau72a]. Zürich [JW08]. zweckmässige [Gau54a].

References

Antosiewicz:1962:NMO

[AG62] Henry A. Antosiewicz and Walter Gautschi. Numerical methods in ordinary differential equations. In Survey of numerical analysis, pages 314–346. McGraw-Hill, New York, 1962.

Anderson:1975:OWC

[AG75] L. A. Anderson and W. Gautschi. Optimal weighted Chebyshevtype quadrature formulas. Calcolo: a quarterly on numerical analysis and theory of computation, 12(3):211–248, 1975. CODEN CALOBK. ISSN 0008-0624 (print), 1126-5434 (electronic).

Alzer:2003:GHM

[Alz03] Horst Alzer. On Gautschi’s harmonic mean inequality for the gamma function. Journal of Computational and Applied Mathematics, 157(1):243–249, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042703004564.

Anonymous:1989:EGK

[Ano89] Anonymous. Erratum: Gauss–Kronrod quadrature formulae for weight functions of Bernstein–Szegő type: Journal of Computational and Applied Mathematics 25 (2) (1989) 199–224. Journal of Computational and Applied Mathematics, 27(3):429, November 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042789900289. See [GN89].

Anonymous:1998:TIC

[Ano98] Anonymous, editor. Tricomi’s ideas and contemporary applied mathematics: (Roma, 28–29 novembre — Torino, 1–2 dicembre 1997), volume 147 of Atti dei convegni Lincei. Accademia Nazionale dei Lincei, Roma, Italy, 1998. ISSN 0391-805X. LCCN QA299.6.

Anonymous:2003:BRIc

[Ano03] Anonymous. Book review: Introduction to numerical analysis, Texts in Applied Mathematics 12: Third Edition. By J. Stoer and R. Bulirsch. Translated by R. Bartels, W. Gautschi, and C. Witzgall. Springer, New York. (2002). 744 pages. $69.95. Computers and Mathematics with Applications, 46(2–3):509–510, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122103900587.

Abramowitz:1964:HMF

[AS64] Milton Abramowitz and Irene A. Stegun, editors. Handbook of Mathematical Functions with Formulas, Graphs, and Mathematical Tables, volume 55 of Applied mathematics series. U. S. Department of Commerce, Washington, DC, USA, 1964. xiv + 1046 pp. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available

online at http://www.convertit.com/Go/ConvertIt/Reference/

AMS55.ASP in bitmap image format.

Askey:1975:TAS

[Ask75] Richard Askey, editor. Theory and application of special functions: proceedings of an advanced seminar sponsored by the Mathematics Research Center, the University of Wisconsin-Madison, March 31–April 2, 1975, number 35 in Publication of the Mathematics Research Center, the University of Wisconsin. Academic Press, New York, NY, USA, 1975. ISBN 0-12-064850-4. LCCN QA3 .U45 no. 35 QA351.

Butzer:1981:BCI

[BF81] Paul Leo Butzer and F. (Franziska) Fehér, editors. E. B. Christoffel, the influence of his work on mathematics and the physical sciences. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1981. ISBN 3-7643-1162-2. LCCN QA7 I57 1981; QA 7 .I57 1981; QA76 .I556 1981. URL https://link.springer.com/

book/10.1007/978-3-0348-5452-8.

Botchev:2006:GTS

[BHvdV06] M. A. Botchev, D. Harutyunyan, and J. J. W. van der Vegt. The Gautschi time stepping scheme for edge finite element discretizations of the Maxwell equations. Journal of Computational Physics, 216(2):654–686, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106000052.

Boisvert:2001:HMF

[BL01] Ronald F. Boisvert and Daniel W. Lozier. Handbook of Mathematical Functions. In Lide [Lid01], pages 135–139.

URL https://nvlpubs.nist.gov/nistpubs/sp958-lide/135-

139.pdf;   https://nvlpubs.nist.gov/nistpubs/sp958-lide/

html/135-139.html. NIST Special Publication.

Brezinski:2013:WGV

[BS13] Claude Brezinski and Ahmed Sameh. Walter Gautschi, Volume 1: Selected Works with Commentaries. Contemporary Mathematicians. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-7033-1, 1-4614-7034-X. xi

+ 694 + 50 pp. LCCN QA297. URL http://site.ebrary.com/
id/10787871.

Brezinski:2014:WGVa

[BS14a] Claude Brezinski and Ahmed Sameh, editors. Walter Gautschi. Volume 2: selected works with commentaries. Contemporary mathematicians. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2014. ISBN 1-4614-7048-X, 1-4614-7049-8 (ebook). xiii + 914 + 33 pp. LCCN QA404.5. URL http://link.

springer.com/10.1007/978-1-4614-7049-6.

Brezinski:2014:WGVb

[BS14b] Claude Brezinski and Ahmed Sameh, editors. Walter Gautschi. Volume 3: selected works with commentaries. Contemporary mathematicians. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2014. ISBN 1-4614-7131-1, 1-4614-7132-X (ebook). xi + 767 + 91 + 29 pp. LCCN QA431. URL http://link.

springer.com/10.1007/978-1-4614-7132-5.

Calio:1986:CGK

[CGM86a] Franca Caliò, Walter Gautschi, and Elena Marchetti. On computing Gauss–Kronrod quadrature formulae. Mathematics of Computation, 47(176):639–650, S57–S63, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Calio:1986:SCG

[CGM86b] Franca Calio, Walter Gautschi, and Elena Marchetti. Supplement to on computing Gauss–Kronrod quadrature formulae. Mathematics of Computation, 47(176):S57–S63, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cordova:1990:VMC

[CGR90] Antonio Córdova, Walter Gautschi, and Stephan Ruscheweyh. Vandermonde matrices on the circle: spectral properties and conditioning. Numerische Mathematik, 57(6/7):577–591, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Davis:2004:IWG

[Dav04] Philip Davis. Interview with Walter Gautschi. SIAM Oral History interview., December 7, 2004. URL https://history.siam.org/

oralhistories/gautschi.htm.

Ellis:1974:RAHb

[Ell74] T. M. R. Ellis. Remark on “Algorithm 420: Hidden-line plotting program”. Communications of the ACM, 17(12):706–??, 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Elliott:1996:BRA

[Ell96] David Elliott. Book reviews: Approximation and Computation: A Festschrift in honor of Walter Gautschi, edited by R. V. M. Zahar. Mathematics of Computation, 65(215):??, July 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00705-3&u=/mcom/1996-65-215/.

Fellmann:2007:LE

[Fel07] Emil Alfred Fellmann. Leonhard Euler. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2007. ISBN 3-7643-7538-8, 3-7643-7539-6 (e-book). xv + 179 pp. LCCN

QA29.E8 F452 2007. URL http://www.loc.gov/catdir/toc/

fy0709/2006937470.html. Translated by E. Gautschi and W. Gautschi from the 1995 German original of the same title.

Flury:1986:ASO

[FG86] Bernhard N. Flury and Walter Gautschi. An algorithm for simultaneous orthogonal transformation of several positive definite symmetric matrices to nearly diagonal form. SIAM Journal on Scientific and Statistical Computing, 7(1):169–184, January 1986. CO-DEN SIJCD4. ISSN 0196-5204.

Frontini:1987:MPS

[FGM87] Marco Frontini, Walter Gautschi, and Gradimir V. Milovanović. Moment-preserving spline approximation on finite intervals. Numerische Mathematik, 50(5):503–518, March 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forster:1982:BOT

[För82] Klaus-Jürgen Förster. Bemerkungen zur optimalen Tschebyscheff-Typ Quadratur. (German) [Remarks on optimal Chebyshev-type quadratures]. Numerische Mathematik, 38(3):421–425, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

G:1968:BRD

[G.68] W. G. Book review: Donald Greenspan, editor, Numerical Solutions of Nonlinear Differential Equations, John Wiley & Sons, Inc., New York, 1966, x + 347 pp. Mathematics of Computation, 22(104):896– 898, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL https://doi.org/10.2307/2004598.

G:1970:BRR

[G.70] W. G. Book review: R. Sauer, I. Szabo, T. P. Angelitch, G. Aumann, F. L. Bauer, R. Bulirsch, H. P. Kunzi, H. Rutishauser, K. Samelson, R. Sauer, and J. Stoer, Mathematische Hilfsmittel des Ingenieurs. III, Springer-Verlag, Berlin, 1969, xviv + 534 pp., 24 cm. Price $24.50. Mathematics of Computation, 24(110):475, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.jstor.org/stable/2004494.

Gautschi:1953:AGI

[Gau53a] Walter Gautschi. Analyse graphischer Integrationsmethoden. (German) [Analysis of graphical integration methods]. Dr. phil., Universität Basel, Basel, Switched, 1953.

Gautschi:1953:ABPa

[Gau53b] Walter Gautschi. The asymptotic behaviour of powers of matrices. Duke Mathematical Journal, 20:127–140, 1953. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Gautschi:1953:ABPb

[Gau53c] Walter Gautschi. The asymptotic behaviour of powers of matrices. II. Duke Mathematical Journal, 20:375–379, 1953. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Gautschi:1953:FGI

[Gau53d] Walter Gautschi. Fehlerabschätzungen für die graphischen Integrationsverfahren von Grammel und Meissner-Ludwig. (German)

[Error estimates for the graphical integration procedures of Grammel and Meissner-Ludwig]. Verhandlungen der Naturforschenden Gesellschaft in Basel, 64:401–435, 1953.

Gautschi:1954:ZUZ

[Gau54a] Walter Gautschi. Über die zeichnerischen Ungenauigkeiten und die zweckmässige Bemessung der Schrittlänge beim graphischen Integrationsverfahren von Meissner-Ludwig. (German) [On the graphical inaccuracies and the appropriate dimensioning of the step length in the graphical integration method by Meissner-Ludwig]. Verhandlungen der Naturforschenden Gesellschaft in Basel, 65:49–66, 1954. CODEN VNGBAH. ISSN 0077-6122.

Gautschi:1954:KLS

[Gau54b] Walter Gautschi. Über eine Klasse von linearen Systemen mit konstanten Koeffizienten. (German) [On a class of linear systems with constant coefficients]. Commentarii Mathematici Helvetici (Zurich), 28:186–196, 1954. CODEN COMHAX. ISSN 0010-2571 (print), 1420-8946 (electronic).

Gautschi:1955:FRK

[Gau55] Walter Gautschi. Über den Fehler des Runge–Kutta-Verfahrens für die numerische Integration gewöhnlicher Differentialgleichungen nter Ordnung. (German) [On the error of the Runge–Kutta method for the numerical integration of ordinary differential equations of n-th order]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 6:456–461, 1955. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Gautschi:1956:BNB

[Gau56a] Walter Gautschi. Bemerkung zu einer notwendigen Bedingung von Picone in der Variationsrechnung. (German) [Remark on a necessary condition of Picone in the variation calculation]. Commentarii Mathematici Helvetici (Zurich), 31:1–4, 1956. CODEN COMHAX. ISSN 0010-2571 (print), 1420-8946 (electronic).

Gautschi:1956:EAI

[Gau56b] Walter Gautschi. Una estensione agli integrali doppi di una condizione di Picone, necessaria per un estremo. (Italian) [An extension to double integrals of a condition of Picone, necessary for one extreme]. Atti della Accademia Nazionale dei Lincei. Rendiconti. Classe di Scienze Fisiche, Matematiche e Naturali. Serie VIII, 20: 283–289, 1956. CODEN ???? ISSN 0392-7881.

Gautschi:1959:EIL

[Gau59a] W. Gautschi. Exponential integral ∫_{1}^{∞} e^{−xt}t^{−n} dt for large values of n. Journal of Research of the National Bureau of Standards (1934), 62(3):123–125, March 1959. ISSN 0091-0635.

Gautschi:1959:NBL

[Gau59b] Walter Gautschi. Note on bivariate linear interpolation for analytic functions. Mathematical Tables and Other Aids to Computation, 13 (66):91–96, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gautschi:1959:SEI

[Gau59c] Walter Gautschi. Some elementary inequalities relating to the gamma and incomplete gamma function. Journal of Mathematics and Physics (MIT), 38(1–4):77–81, April 1959. CODEN JMPHA9.

ISSN 0097-1421. URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm195938177.

Gautschi:1961:NIO

[Gau61a] Walter Gautschi. Numerical integration of ordinary differential equations based on trigonometric polynomials. Numerische Mathematik, 3(1):381–397, December 1961. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=3&issue=1&spage=381.

Gautschi:1961:RCC

[Gau61b] Walter Gautschi. Recursive computation of certain integrals. Journal of the ACM, 8(1):21–40, January 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735x (electronic).

Gautschi:1961:RCR

[Gau61c] Walter Gautschi. Recursive computation of the repeated integrals of the error function. Mathematics of Computation, 15(75):227–232, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1962:DFS

[Gau62a] Walter Gautschi. Diffusion functions for small argument. SIAM Review, 4(3):227–229, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gautschi:1962:ILS

[Gau62b] Walter Gautschi. Instability of linear second-order difference equations [abstract only]. In Proceedings of the International Conference on Information Processing, page 207. ????, ????, 1962.

Gautschi:1962:IVC

[Gau62c] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices. Numerische Mathematik, 4:117–123, 1962. CO-DEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Gautschi:1963:IVC

[Gau63a] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices. II. Numerische Mathematik, 5:425–430, 1963. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Gautschi:1963:RCS

[Gau63b] Walter Gautschi. Recursive computation of special functions. In ????, editor, The University of Michigan Engineering Summer Conferences, Numerical Analysis, Summer 1963, page ?? ????, ????, 1963.

Gautschi:1964:AAI

[Gau64a] W. Gautschi. ACM Algorithm 222: Incomplete beta function ratios. Communications of the ACM, 7(3):143–144, March 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [KD71, Ell74].

Gautschi:1964:AAB

[Gau64b] W. Gautschi. ACM Algorithm 236: Bessel functions of the first kind [S17]. Communications of the ACM, 7(8):479–480, August 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sko75].

Gautschi:1964:AGF

[Gau64c] Walter Gautschi. Algorithm 221: Gamma functions. Communications of the ACM, 7(3):143, March 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1964:AIB

[Gau64d] Walter Gautschi. Algorithm 222: Incomplete beta functions ratios. Communications of the ACM, 7(3):143, March 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [KD71, Ell74].

Gautschi:1964:CAI

[Gau64e] Walter Gautschi. Certification of Algorithm 222: Incomplete beta function ratios. Communications of the ACM, 7(4):244, April 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1964:EFF

[Gau64f] Walter Gautschi. Error function and Fresnel integrals. In Abramowitz and Stegun [AS64], pages 295–330. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://www.

convertit.com/Go/ConvertIt/Reference/AMS55.ASP in bitmap

image format.

Gautschi:1965:ALF

[Gau65a] W. Gautschi. Algorithm 259: Legendre functions for arguments larger than one [S16]. Communications of the ACM, 8(8):488–492, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Jan77].

Gautschi:1965:CAS

[Gau65b] Walter Gautschi. Certification of Algorithm 236 [S17]: Bessel functions of the first kind. Communications of the ACM, 8(2):105–106, February 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1966:AD

[Gau66a] Walter Gautschi. Algorithm 282: Derivatives of e^{x}/x, cos(x)/x, and sin(x)/x. Communications of the ACM, 9(4):272, April 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [GK70b].

Gautschi:1966:ARC

[Gau66b] Walter Gautschi. Algorithm 292: Regular Coulomb wave functions. Communications of the ACM, 9(11):793–795, November 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1966:CSD

[Gau66c] Walter Gautschi. Computation of successive derivatives of f(z)/z. Mathematics of Computation, 20(94):209–214, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1967:CAT

[Gau67a] Walter Gautschi. Computational aspects of three-term recurrence relations. SIAM Review, 9(1):24–82, January 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://link.aip.org/link/?SIR/9/24/1.

Gautschi:1967:NQP

[Gau67b] Walter Gautschi. Numerical quadrature in the presence of a singularity. SIAM Journal on Numerical Analysis, 4(3):357–362, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gautschi:1968:AAG

[Gau68a] Walter Gautschi. Algorithm 331: Gaussian quadrature formulas [D1]. Communications of the ACM, 11(6):432–436, June 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1968:CGC

[Gau68b] Walter Gautschi. Construction of Gauss–Christoffel quadrature formulas. Mathematics of Computation, 22(102):251–270, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1969:AAC

[Gau69a] W. Gautschi. ACM Algorithm 363: Complex error function. Communications of the ACM, 12(11):635, November 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Köl72].

Gautschi:1969:ACE

[Gau69b] Walter Gautschi. Algorithm 363: Complex error function [S15]. Communications of the ACM, 12(11):635, November 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Köl72].

Gautschi:1969:AMS

[Gau69c] Walter Gautschi. An application of minimal solutions of three-term recurrences to Coulomb wave functions. Aequationes Mathematicae, 2:171–176, 1969. CODEN AEMABN. ISSN 0001-9054 (print), 1420-8903 (electronic).

Gautschi:1969:CMA

[Gau69d] Walter Gautschi. On the condition of a matrix arising in the numerical inversion of the Laplace transform. Mathematics of Computation, 23(105):109–118, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1969:RAS

[Gau69e] Walter Gautschi. Remark on Algorithm 292 [S22]: Regular Coulomb wave functions. Communications of the ACM, 12(5):280, May 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1970:ECC

[Gau70a] Walter Gautschi. Efficient computation of the complex error function. SIAM Journal on Numerical Analysis, 7(1):187–198, March 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [Wei95].

Gautschi:1970:CGQ

[Gau70b] Walter Gautschi. On the construction of Gaussian quadrature rules from modified moments. Mathematics of Computation, 24(110): 245–260, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1970:TGQ

[Gau70c] Walter Gautschi. Tables of Gaussian quadrature rules for the calculation of Fourier coefficients. Mathematics of Computation, 24(110, loose microfiche suppl):A–D, 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1972:NRR

[Gau72a] W. Gautschi. Zur Numerik rekurrenter Relationen. (German) [Numerical aspects of recurrence relations]. Computing: Archiv für Informatik und Numerik, 9(2):107–126, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gautschi:1972:AFP

[Gau72b] Walter Gautschi. Attenuation factors in practical Fourier analysis. Numerische Mathematik, 18(5):373–400, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1972:COP

[Gau72c] Walter Gautschi. The condition of orthogonal polynomials. Mathematics of Computation, 26(120):923–924, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1973:AAE

[Gau73a] Walter Gautschi. ACM Algorithm 471: Exponential integrals [S13]. Communications of the ACM, 16(12):761–763, December 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1973:NAR

[Gau73b] Walter Gautschi. Numerical aspects of recurrence relations. Aerospace Research Laboratories, Air Force Systems Command, United States Air Force, Wright-Patterson Air Force Base, Ohio, 1973. v + 38 pp. Translated by the author, with the publisher’s permission, from Computing (Arch. Elektron. Rechnen) 9 (1972), 107–126, ARL 73-0005.

Gautschi:1973:CAE

[Gau73c] Walter Gautschi. On the condition of algebraic equations. Numerische Mathematik, 21(5):405–424, October 1973. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Gautschi:1974:HMI

[Gau74a] Walter Gautschi. A harmonic mean inequality for the gamma function. SIAM Journal on Mathematical Analysis, 5(2):278–281, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gautschi:1974:SMV

[Gau74b] Walter Gautschi. Some mean value inequalities for the gamma function. SIAM Journal on Mathematical Analysis, 5(2):282–292, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gautschi:1975:CMS

[Gau75a] Walter Gautschi. Computational methods in special functions — a survey. In Askey [Ask75], pages 1–98. Math. Res. Center, Univ. Wisconsin Publ., No. 35. ISBN 0-12-064850-4. LCCN QA3 .U45 no. 35 QA351.

Gautschi:1975:NCT

[Gau75b] Walter Gautschi. Nonexistence of Chebyshev-type quadratures on infinite intervals. Mathematics of Computation, 29(129):93–99, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1975:NEI

[Gau75c] Walter Gautschi. Norm estimates for inverses of Vandermonde matrices. Numerische Mathematik, 23(4):337–347, August 1975. CO-DEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Gautschi:1975:OCV

[Gau75d] Walter Gautschi. Optimally conditioned Vandermonde matrices. Numerische Mathematik, 24(1):1–12, February 1975. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Gautschi:1975:SDG

[Gau75e] Walter Gautschi. Stime dell’errore globale nei metodi “one-step” per equazioni differenziali ordinarie. (Italian) [Estimates of the global error in “one-step” methods for ordinary differential equations]. Rendiconti di Matematica. Serie VII. Università degli Studi di Roma. Istituto Matematica Guido Castelnuovo. Istituto di Matematica Applicata. Istituto Nazionale di Alta Matematica, 8(2): 601–617, 1975. ISSN 0034-4427. Collection of articles dedicated to Mauro Picone, II.

Gautschi:1976:ACQ

[Gau76] Walter Gautschi. Advances in Chebyshev quadrature. Lecture Notes in Mathematics, 506:100–121, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.com/

chapter/10.1007/BFb0080118/.

Gautschi:1977:ARI

[Gau77a] Walter Gautschi. Algorithm 521: Repeated integrals of the coerror function [S15]. ACM Transactions on Mathematical Software, 3(3): 301–302, September 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gautschi:1977:ACC

[Gau77b] Walter Gautschi. Anomalous convergence of a continued fraction for ratios of Kummer functions. Mathematics of Computation, 31 (140):994–999, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1977:ERI

[Gau77c] Walter Gautschi. Evaluation of repeated integrals of the coerror function. ACM Transactions on Mathematical Software, 3(3):240– 252, September 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gautschi:1978:IVC

[Gau78a] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices. III. Numerische Mathematik, 29(4):445–450, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1978:QNC

[Gau78b] Walter Gautschi. Questions of numerical conditions related to polynomials. In Recent advances in numerical analysis (Proc. Sympos., Math. Res. Center, Univ. Wisconsin, Madison, Wis., 1978), volume 41 of Publ. Math. Res. Center Univ. Wisconsin, pages 45–72. Academic Press, New York, NY, USA, 1978.

Gautschi:1979:AIG

[Gau79a] W. Gautschi. Algorithm 542: Incomplete gamma functions [S14]. ACM Transactions on Mathematical Software, 5(4):482–489, December 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gautschi:1979:CPI

[Gau79b] Walter Gautschi. A computational procedure for incomplete gamma functions. ACM Transactions on Mathematical Software, 5(4):466– 481, December 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gautschi:1979:CPP

[Gau79c] Walter Gautschi. The condition of polynomials in power form. Mathematics of Computation, 33(145):343–352, January 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1979:PPL

[Gau79d] Walter Gautschi. On the preceding paper “A Legendre polynomial integral” by James L. Blue. Mathematics of Computation, 33(146):

742–743, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1981:CNA

[Gau81a] Walter Gautschi. Christoffel and numerical analysis. In Butzer and Fehér [BF81], page 749. ISBN 3-7643-1162-2. LCCN QA7 I57 1981; QA 7 .I57 1981; QA76 .I556 1981. URL https://link.springer.

com/book/10.1007/978-3-0348-5452-8.

Gautschi:1981:MST

[Gau81b] Walter Gautschi. Minimal solutions of three-term recurrence relations and orthogonal polynomials. Mathematics of Computation, 36(154):547–554, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1981:RCW

[Gau81c] Walter Gautschi. Recognition of Christoffel’s work on quadrature during and after his lifetime. In Butzer and Fehér [BF81], pages 724–727. ISBN 3-7643-1162-2. LCCN QA7 I57 1981; QA 7 .I57

1981; QA76 .I556 1981. URL https://link.springer.com/book/
10.1007/978-3-0348-5452-8.

Gautschi:1981:SGC

[Gau81d] Walter Gautschi. A survey of Gauss–Christoffel quadrature formulae. In Butzer and Fehér [BF81], pages 72–147. ISBN 3-7643-1162-2. LCCN QA7 I57 1981; QA 7 .I57 1981; QA76 .I556

1981. URL https://link.springer.com/book/10.1007/978-3-
0348-5452-8.

Gautschi:1982:GOP

[Gau82a] Walter Gautschi. On generating orthogonal polynomials. SIAM Journal on Scientific and Statistical Computing, 3(3):289–317, September 1982. CODEN SIJCD4. ISSN 0196-5204.

Gautschi:1982:POR

[Gau82b] Walter Gautschi. Polynomials orthogonal with respect to the reciprocal gamma function. BIT (Nordisk tidskrift for informationsbehandling), 22(3):387–389, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=22&issue=3&spage=387.

Gautschi:1983:CVL

[Gau83a] Walter Gautschi. The condition of Vandermonde-like matrices involving orthogonal polynomials. Linear Algebra and its Applications, 52/53(??):293–300, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gautschi:1983:HHC

[Gau83b] Walter Gautschi. How and how not to check Gaussian quadrature formulae. BIT (Nordisk tidskrift for informationsbehandling), 23(2): 209–216, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=23&
issue=2&spage=209.

Gautschi:1983:CBC

[Gau83c] Walter Gautschi. On the convergence behavior of continued fractions with real elements. Mathematics of Computation, 40(161): 337–342, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1983:AMO

[Gau83d] Walter Gautschi. To Alexander M. Ostrowski on his ninetieth birthday. Linear Algebra and its Applications, 52/53(??):xi–xiv, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gautschi:1984:DAS

[Gau84a] Walter Gautschi. Discrete approximations to spherically symmetric distributions. Numerische Mathematik, 44(1):53–60, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1984:SOP

[Gau84b] Walter Gautschi. On some orthogonal polynomials of interest in theoretical chemistry. BIT (Nordisk tidskrift for informationsbehandling), 24(4):473–483, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=24&issue=4&spage=473.

Gautschi:1984:QNC

[Gau84c] Walter Gautschi. Questions of numerical condition related to polynomials. In Golub [Gol84], pages 140–177. ISBN 0-88385-126-1 (v. 1), 0-88385-100-8 (set). LCCN QA297 .S83 1984.

Gautschi:1985:OPC

[Gau85a] Walter Gautschi. Orthogonal polynomials — constructive theory and applications. Journal of Computational and Applied Mathematics, 12–13(??):61–76, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037704278590007X.

Gautschi:1985:SNA

[Gau85b] Walter Gautschi. Some new applications of orthogonal polynomials. Lecture Notes in Mathematics, 1171:63–73, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076531/.

Gautschi:1986:SOP

[Gau86] Walter Gautschi. On the sensitivity of orthogonal polynomials to perturbations in the moments. Numerische Mathematik, 48(4):369– 382, April 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1987:CIH

[Gau87] Walter Gautschi. A conjectured inequality for Hermite interpolation at the zeros of Jacobi polynomials. SIAM Review, 29(2):297–298, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gautschi:1989:ZPO

[Gau89] Walter Gautschi. On the zeros of polynomials orthogonal on the semicircle. SIAM Journal on Mathematical Analysis, 20(3):738–743, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gautschi:1990:HSV

[Gau90] Walter Gautschi. How (un)stable are Vandermonde systems? In Wong [Won90], pages 193–210. ISBN 0-8247-8347-6. LCCN

QA299.6 .A88 1990. URL http://www.loc.gov/catdir/enhancements/
fy0647/90002810-d.html.

Gautschi:1991:CSCa

[Gau91a] Walter Gautschi. A class of slowly convergent series and their summation by Gaussian quadrature. Mathematics of Computation, 57 (195):309–324, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1991:CSCb

[Gau91b] Walter Gautschi. On certain slowly convergent series occurring in plate contact problems. Mathematics of Computation, 57(195):325– 338, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1991:QFH

[Gau91c] Walter Gautschi. Quadrature formulae on half-infinite intervals. BIT (Nordisk tidskrift for informationsbehandling), 31(3):437–446, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=31&
issue=3&spage=437.

Gautschi:1992:MCE

[Gau92] Walter Gautschi. On mean convergence of extended Lagrange interpolation. Journal of Computational and Applied Mathematics, 43(1–2):19–35, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037704279290257X.

Gautschi:1993:RRO

[Gau93a] Walter Gautschi. Is the recurrence relation for orthogonal polynomials always stable? BIT (Nordisk tidskrift for informationsbehandling), 33(2):277–284, June 1993. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.se/BIT/contents/bit33.
html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=33&issue=2&spage=277.

Gautschi:1993:CGF

[Gau93b] Walter Gautschi. On the computation of generalized Fermi–Dirac and Bose–Einstein integrals. Computer Physics Communications, 74(2):233–238, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/001046559390093R.

Gautschi:1994:ACP

[Gau94a] Walter Gautschi. Algorithm 726: ORTHPOL—a package of routines for generating orthogonal polynomials and Gauss-type quadrature rules. ACM Transactions on Mathematical Software, 20 (1):21–62, March 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/pubs/

citations/journals/toms/1994-20-1/p21-gautschi/. See re-

mark [Gau98b].

Gautschi:1994:MCH

[Gau94b] Walter Gautschi, editor. Mathematics of computation, 1943–1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gautschi:1995:WPR

[Gau95] Walter Gautschi. The work of Philip Rabinowitz on numerical integration. Numerical Algorithms, 9(3–4):199–222, ???? 1995. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gautschi:1996:BRP

[Gau96a] Walter Gautschi. Book reviews: Parallel processing for scientific computing, edited by David H. Bailey, Petter E. Bjørstad, John R. Gilbert, Michael V. Mascagni, Robert S. Schreiber, Horst D. Simon, Virginia J. Torczon and Layne T. Watson. Mathematics of Computation, 65(213):??, January 1996. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-7&u=/mcom/1996-65-213/.

Gautschi:1996:OPA

[Gau96b] Walter Gautschi. Orthogonal polynomials: applications and computation. Acta Numerica, 5:45–119, 1996. CODEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Gautschi:1997:MQP

[Gau97a] W. Gautschi. Moments in quadrature problems. Computers and Mathematics with Applications, 33(1–2):105–118, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122196002234.

Gautschi:1997:NA

[Gau97b] Walter Gautschi. Numerical analysis. Birkhäuser Boston Inc., Cambridge, MA, USA, 1997. ISBN 0-8176-3895-4. xiv + 506 pp. An introduction.

Gautschi:1997:NAI

[Gau97c] Walter Gautschi. Numerical analysis: an introduction. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1997. ISBN 0-8176-3895-4, 3-7643-3895-4 (Basel). xiii + 506 pp. LCCN QA297 .G35 1997.

Gautschi:1998:IGF

[Gau98a] Walter Gautschi. The incomplete gamma functions since Tricomi. In Anonymous [Ano98], pages 203–237. ISSN 0391-

805X. LCCN QA299.6. URL https://web.archive.org/web/
20070503103643/http://citeseer.ist.psu.edu/gautschi98incomplete.
html.

Gautschi:1998:RAO

[Gau98b] Walter Gautschi. Remark on Algorithm 726: ORTHPOL — a package of routines for generating orthogonal polynomials and Gauss-type quadrature rules. ACM Transactions on Mathematical Software, 24(3):355, September 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org:80/pubs/citations/journals/toms/1998-24-3/p355-

gautschi/. See [Gau94a].

Gautschi:1999:AGG

[Gau99a] Walter Gautschi. Algorithm 793: GQRAT — Gauss quadrature for rational functions. ACM Transactions on Mathematical Software, 25(2):213–239, June 1999. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

ftp://netlib.bell-labs.com/netlib/toms/793.gz;   http:/
/phase.etl.go.jp/netlib/toms/793;   http://www.acm.org/
pubs/citations/journals/toms/1999-25-2/p213-gautschi/;
 http://www.hensa.ac.uk/netlib/toms/793.gz;   http://www.
netlib.no/netlib/toms/793; http://www.netlib.org/toms/

Gautschi:1999:NRC

[Gau99b] Walter Gautschi. A note on the recursive calculation of incomplete gamma functions. ACM Transactions on Mathematical Software, 25(1):101–107, March 1999. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org:80/

pubs/citations/journals/toms/1999-25-1/p101-gautschi/.

Gautschi:1999:OPQ

[Gau99c] Walter Gautschi. Orthogonal polynomials and quadrature. Electronic Transactions on Numerical Analysis, 9:65–76, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.9.1999/pp65-76.dir/pp65-76.pdf.

Gautschi:2000:BRC

[Gau00a] Walter Gautschi. Book review: Computational Integration. Mathematics of Computation, 69(231):1311–1312, July 2000. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/mcom/2000-69-231/S0025-5718-00-
01249-7/bookrev-S0025-5718-00-01249-7.html;   http://www.
ams.org/mcom/2000-69-231/S0025-5718-00-01249-7/S0025-5718-
00-01249-7.dvi;   http://www.ams.org/mcom/2000-69-231/
S0025-5718-00-01249-7/S0025-5718-00-01249-7.pdf;   http:
//www.ams.org/mcom/2000-69-231/S0025-5718-00-01249-7/S0025-
5718-00-01249-7.ps;   http://www.ams.org/mcom/2000-69-
231/S0025-5718-00-01249-7/S0025-5718-00-01249-7.tex.

Gautschi:2000:GRF

[Gau00b] Walter Gautschi. Gauss–Radau formulae for Jacobi and Laguerre weight functions. Mathematics and Computers in Simulation, 54 (4–5):403–412, December 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0378475400001798.

Gautschi:2000:HOG

[Gau00c] Walter Gautschi. High-order Gauss–Lobatto formulae. Numerical Algorithms, 25(1–4):213–222, September 2000. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/5058/
3/20/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/3/20/fulltext.pdf. Mathematical jour-

ney through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Gautschi:2001:NBF

[Gau01a] Walter Gautschi. Note: Barycentric formulae for cardinal (SINC- ) interpolants by Jean-Paul Berrut. Numerische Mathematik, 87 (4):791–792, February 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1087004/10870791.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1087004/10870791.pdf.

Gautschi:2001:URF

[Gau01b] Walter Gautschi. The use of rational functions in numerical quadrature. Journal of Computational and Applied Mathematics, 133(1– 2):111–126, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0377042700006373.

Gautschi:2002:GQA

[Gau02a] W. Gautschi. Gauss quadrature approximations to hypergeometric and confluent hypergeometric functions. Journal of Computational and Applied Mathematics, 139(1):173–187, February 1, 2002. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Gautschi:2002:CBA

[Gau02b] Walter Gautschi. Computation of Bessel and Airy functions and of related Gaussian quadrature formulae. BIT Numerical Mathematics, 42(1):110–118, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=42&issue=1&spage=110.

Gautschi:2002:IBC

[Gau02c] Walter Gautschi. The interplay between classical analysis and (numerical) linear algebra — a tribute to Gene H. Golub. Electronic Transactions on Numerical Analysis, 13(??):119–147, ???? 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

URL http://etna.mcs.kent.edu/vol.13.2002/pp119-147.dir/
pp119-147.pdf.

Gautschi:2004:GGR

[Gau04a] Walter Gautschi. Generalized Gauss–Radau and Gauss–Lobatto formulae. BIT Numerical Mathematics, 44(4):711–720, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=44&
issue=4&spage=711.

Gautschi:2004:OPC

[Gau04b] Walter Gautschi. Orthogonal Polynomials: Computation and Approximation. Oxford University Press, Oxford, UK, 2004. ISBN 0-19-850672-4. viii + 301 pp. LCCN QA404.5 .G356 2004.

Gautschi:2005:CPO

[Gau05a] Walter Gautschi. Computing polynomials orthogonal with respect to densely oscillating and exponentially decaying weight functions and related integrals. Journal of Computational and Applied Mathematics, 184(2):493–504, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042705000440.

Gautschi:2005:HLF

[Gau05b] Walter Gautschi. The Hardy–Littlewood function: an exercise in slowly convergent series. Journal of Computational and Applied Mathematics, 179(1–2):249–254, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042704004510.

Gautschi:2005:HNG

[Gau05c] Walter Gautschi. A historical note on Gauss–Kronrod quadrature. Numerische Mathematik, 100(3):483–484, May 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:2005:NQC

[Gau05d] Walter Gautschi. Numerical quadrature computation of the Macdonald function for complex orders. BIT Numerical Mathematics, 45(3):593–603, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=45&issue=3&spage=593.

Gautschi:2005:OPM

[Gau05e] Walter Gautschi. Orthogonal polynomials (in Matlab). Journal of Computational and Applied Mathematics, 178(1–2):215– 234, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042704003760.

Gautschi:2006:CTR

[Gau06a] Walter Gautschi. The circle theorem and related theorems for Gauss-type quadrature rules. Electronic Transactions on Numerical Analysis, 25:129–137, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://etna.mcs.kent.edu/

vol.25.2006/pp129-137.dir/pp129-137.pdf.

Gautschi:2006:CKL

[Gau06b] Walter Gautschi. Computing the Kontorovich–Lebedev integral transforms and their inverses. BIT Numerical Mathematics, 46(1): 21–40, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gautschi:2006:OPQ

[Gau06c] Walter Gautschi. Orthogonal polynomials, quadrature, and approximation: Computational methods and software (in Matlab). Lecture Notes in Mathematics, 1883:1–77, 2006. CO-DEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/content/pdf/10.1007/978-3-540-
36716-1_1.pdf;   http://link.springer.com/content/pdf/
bfm:978-3-540-36716-1/1.pdf.

Gautschi:2007:GTT

[Gau07] Walter Gautschi. A guided tour through my bibliography. Numerical Algorithms, 45(1–4):11–35, August 2007. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=45&issue=1&spage=11.

Gautschi:2008:ESL

[Gau08a] Walter Gautschi. Euler special lecture: Leonhard Euler: His life, the man, and his works. In Jeltsch and Wanner [JW08], pages 447– ?? ISBN 3-03719-056-6, 3-03719-556-8 (e-book). LCCN QA1 .I73 2007.

Gautschi:2008:LEH

[Gau08b] Walter Gautschi. Leonhard Euler: His life, the man, and his works. SIAM Review, 50(1):3–33, 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://link.aip.org/link/

?SIR/50/3/1.

Gautschi:2008:NEC

[Gau08c] Walter Gautschi. The numerical evaluation of a challenging integral. Numerical Algorithms, 49(1–4):187–194, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=49&issue=1&spage=187.

Gautschi:2008:CIL

[Gau08d] Walter Gautschi. On a conjectured inequality for the largest zero of Jacobi polynomials. Numerical Algorithms, 49(1–4):195– 198, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=49&
issue=1&spage=195.

Gautschi:2008:EAC

[Gau08e] Walter Gautschi. On Euler’s attempt to compute logarithms by interpolation: a commentary to his letter of February 16, 1734 to Daniel Bernoulli. Journal of Computational and Applied Mathematics, 219(2):408–415, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037704270600714X.

Gautschi:2009:HOG

[Gau09a] Walter Gautschi. High-order generalized Gauss–Radau and Gauss– Lobatto formulae for Jacobi and Laguerre weight functions. Numerical Algorithms, 51(2):143–149, June 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=51&issue=2&spage=143. Tributes to Gene H.

Golub, Part II.

Gautschi:2009:NCI

[Gau09b] Walter Gautschi. New conjectured inequalities for zeros of Jacobi polynomials. Numerical Algorithms, 50(3):293–296, March 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=1017-1398&volume=50&issue=3&spage=293.

See remark [?].

Gautschi:2009:CIZ

[Gau09c] Walter Gautschi. On conjectured inequalities for zeros of Jacobi polynomials. Numerical Algorithms, 50(1):93–96, January 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=1017-1398&volume=50&issue=1&spage=93.

Gautschi:2009:VPR

[Gau09d] Walter Gautschi. Variable-precision recurrence coefficients for nonstandard orthogonal polynomials. Numerical Algorithms, 52(3): 409–418, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=52&
issue=3&spage=409.

Gautschi:2009:HSB

[Gau10a] Walter Gautschi. How sharp is Bernstein’s inequality for Jacobi polynomials? Electronic Transactions on Numerical Analysis, 36: 1–8, 2009–2010. CODEN ???? ISSN 1068-9613 (print), 1097-

4067 (electronic). URL http://etna.mcs.kent.edu/vol.36.
2009-2010/pp1-8.dir/pp1-8.pdf;   http://etna.mcs.kent.
edu/volumes/2001-2010/vol36/abstract.php?vol=36&pages=
1-8.

Gautschi:2010:GQR

[Gau10b] Walter Gautschi. Gauss quadrature routines for two classes of logarithmic weight functions. Numerical Algorithms, 55(2–3): 265–277, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=55&
issue=2&spage=265.

Gautschi:2010:STN

[Gau10c] Walter Gautschi. The spiral of Theodorus, numerical analysis, and special functions. Journal of Computational and Applied Mathematics, 235(4):1042–1052, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042709008012.

Gautschi:2011:EMI

[Gau11a] Walter Gautschi. Experimental mathematics involving orthogonal polynomials. In Gautschi et al. [GMR11], pages 117–134. ISBN 1-4419-6593-9 (paperback), 1-4419-6594-7 (e-book), 1-4419-6595-5, 1-4614-2703-7. LCCN QA39.2 .A67 2011; QA221 .A6345

2011. URL https://link.springer.com/chapter/10.1007/978-
1-4419-6594-3_9.

Gautschi:2011:LWF

[Gau11b] Walter Gautschi. The Lambert W -functions and some of their integrals: a case study of high-precision computation. Numerical Algorithms, 57(1):27–34, May 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=57&
issue=1&spage=27.

Gautschi:2011:MCG

[Gau11c] Walter Gautschi. My collaboration with Gradimir V. Milovanović. In Gautschi et al. [GMR11], pages 33–43. ISBN 1-4419-6593-9 (paperback), 1-4419-6594-7 (e-book), 1-4419-6595-5, 1-4614-2703-7. LCCN QA39.2 .A67 2011; QA221 .A6345 2011.

Gautschi:2011:OSO

[Gau11d] Walter Gautschi. Optimally scaled and optimally conditioned Vandermonde and Vandermonde-like matrices. BIT Numerical Mathematics, 51(1):103–125, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=51&issue=1&spage=103.

Gautschi:2011:RNC

[Gau11e] Walter Gautschi. Remark on “New conjectured inequalities for zeros of Jacobi polynomials” by Walter Gautschi, Numer. Algorithms 50:293–296 (2009). Numerical Algorithms, 57 (4):511, August 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=57&
issue=4&spage=511. See [Gau09b].

Gautschi:2012:IBA

[Gau12a] W. Gautschi. Interpolation before and after Lagrange. Rend. Semin. Mat. Univ. Politec. Torino, 70(4):347–368, 2012. ISSN 0373-1243.

Gautschi:2012:NIS

[Gau12b] Walter Gautschi. Numerical integration over the square in the presence of algebraic/logarithmic singularities with an application to aerodynamics. Numerical Algorithms, 61(2):275– 290, October 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=61&

issue=2&spage=275. See erratum [Gau13a].

Gautschi:2012:SRJ

[Gau12c] Walter Gautschi. Sub-range Jacobi polynomials. Numerical Algorithms, 61(4):649–657, December 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/article/10.1007/s11075-012-9556-z/;
 http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=61&issue=4&spage=649-657. See erra-

tum [Gau17a].

Gautschi:2013:ENI

[Gau13a] Walter Gautschi. Erratum to: Numerical integration over the square in the presence of algebraic/logarithmic singularities with an application to aerodynamics. Numerical Algorithms, 64(4):759, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://link.springer.com/article/
10.1007/s11075-013-9787-7;   http://link.springer.com/
content/pdf/10.1007/s11075-013-9787-7.pdf. See [Gau12b].

Gautschi:2013:NNS

[Gau13b] Walter Gautschi. Neutralizing nearby singularities in numerical quadrature. Numerical Algorithms, 64(3):417–425, November 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9672-9.

Gautschi:2013:RMO

[Gau13c] Walter Gautschi. Repeated modifications of orthogonal polynomials by linear divisors. Numerical Algorithms, 63(2):369–383, June 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9627-1.

Gautschi:2014:HPG

[Gau14] Walter Gautschi. High-precision Gauss–Turán quadrature rules for Laguerre and Hermite weight functions. Numerical Algorithms, 67 (1):59–72, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.com/

article/10.1007/s11075-013-9774-z.

Gautschi:2015:POR

[Gau15] Walter Gautschi. Polynomials orthogonal with respect to exponential integrals. Numerical Algorithms, 70(1):215–226, September 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9943-8.

Gautschi:2016:AER

[Gau16a] Walter Gautschi. Algorithm 957: Evaluation of the repeated integral of the coerror function by half-range Gauss–Hermite quadrature. ACM Transactions on Mathematical Software, 42(1):9:1–9:10, February 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gautschi:2016:OPM

[Gau16b] Walter Gautschi. Orthogonal polynomials in MATLAB: exercises and solutions, volume 26 of Software, environments, and tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2016. ISBN 1-61197-429-1 (paperback), 1-61197-430-5. ix + 335 pp. LCCN QA404.5 .G3564 2016.

Gautschi:2017:ESR

[Gau17a] Walter Gautschi. Erratum to: Sub-range Jacobi polynomials. Numerical Algorithms, 74(2):637, February 2017. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s11075-016-0257-

x.pdf. See [Gau12c].

Gautschi:2017:MPZ

[Gau17b] Walter Gautschi. Monotonicity properties of the zeros of Freud and sub-range Freud polynomials: Analytic and empirical results. Mathematics of Computation, 86(304):855–864, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://www.ams.org/journals/mcom/
2017-86-304/S0025-5718-2016-03181-6;   http://www.ams.
org/journals/mcom/2017-86-304/S0025-5718-2016-03181-6/
S0025-5718-2016-03181-6.pdf;   http://www.ams.org/mathscinet/
search/author.html?mrauthid=71975.

Gautschi:2017:POR

[Gau17c] Walter Gautschi. Polynomials orthogonal with respect to cardinal B-spline weight functions. Numerical Algorithms, 76(4):1099–1107, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gautschi:2018:DTM

[Gau18a] Walter Gautschi. A discrete top-down Markov problem in approximation theory. In Frontiers in orthogonal polynomials and q-series, volume 1 of Contemp. Math. Appl. Monogr. Expo. Lect. Notes, pages 267–289. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2018.

Gautschi:2018:ILA

[Gau18b] Walter Gautschi. On the Ismail–Letessier–Askey monotonicity conjecture for zeros of ultraspherical polynomials. In Frontiers in orthogonal polynomials and q-series, volume 1 of Contemp. Math. Appl. Monogr. Expo. Lect. Notes, pages 251–266. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2018.

Gautschi:2018:ZSJ

[Gau18c] Walter Gautschi. On the zeros of subrange Jacobi polynomials. Numerical Algorithms, 79(3):759–768, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See correction [Gau19].

Gautschi:2018:PAS

[Gau18d] Walter Gautschi. Progress by accident: some reflections on my career. SIAM News, 51(10):7–8, 2018. ISSN 1557-9573.

Gautschi:2018:SRO

[Gau18e] Walter Gautschi. A software repository for orthogonal polynomials, volume 28 of Software, Environments, and Tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2018. ISBN 1-61197-522-0. viii + 60 pp. LCCN QA404.5 .G3567 2018.

Gautschi:2019:CSR

[Gau19] Walter Gautschi. Correction to: Sub-range Jacobi polynomials. Numerical Algorithms, 81(2):771, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11075-019-00715-9.pdf.

See [Gau18c].

Gautschi:2021:SRG

[Gau21] Walter Gautschi. A software repository for Gaussian quadratures and Christoffel functions, volume 32 of Software, Environments, and Tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2021. ISBN 1-61197-634-0 (paperback), 1-61197-635-9 (e-book). xvii + 134 pp. LCCN QA299.4.G3 G38 2021.

Gautschi:2022:ALP

[Gau22] Walter Gautschi. Another look at polynomials orthogonal relative to exponential integral weight functions. Numerical Algorithms, 91(4):1547–1557, December 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01313-y.

Gautschi:1971:AFP

[Gau72] Walter Gautschi. Attenuation factors in practical Fourier analysis. Numerische Mathematik, 18:373–400, 1971/72. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Gautschi:1964:EIR

[GC64] Walter Gautschi and William F. Cahill. Exponential integral and related functions. In Abramowitz and Stegun [AS64], pages 227– 252. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available

online at http://www.convertit.com/Go/ConvertIt/Reference/

AMS55.ASP in bitmap image format.

Gander:1998:AQR

[GG98] Walter Gander and Walter Gautschi. Adaptive quadrature: Revisited. Technical Report 306, ETH Zürich, Zürich, Switzerland, 1998.

ii + 21 pp. URL http://hdl.handle.net/20.500.11850/69327.

Gander:2000:AQR

[GG00] Walter Gander and Walter Gautschi. Adaptive quadrature — revisited. BIT Numerical Mathematics, 40(1):84–101, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125

(electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=40&issue=1&spage=84.

Gautschi:2008:LGW

[GG08] Walter Gautschi and Carla Giordano. Luigi Gatteschi’s work on asymptotics of special functions and their zeros. Numerical Algorithms, 49(1–4):11–31, December 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=49&issue=1&spage=11.

Gautschi:2000:QRR

[GGL00] Walter Gautschi, Laura Gori, and M. Laura Lo Cascio. Quadrature rules for rational functions. Numerische Mathematik, 86(4):617– 633, October 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086004/00860617.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086004/00860617.pdf.

Gautschi:1999:ACO

[GGO99] Walter Gautschi, Gene H. Golub, and Gerhard Opfer, editors. Applications and computation of orthogonal polynomials: conference at the Mathematical Research Institute Oberwolfach, Germany, March 22–28, 1998, volume 131 of International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 3-7643-6137-9, 0-8176-6137-9 (paperback). LCCN QA404.5 .A67 1999.

Gautschi:2019:CST

[GH19] Walter Gautschi and Ernst Hairer. On conjectures of Stenger in the theory of orthogonal polynomials. Journal of Inequalities and Applications, page 27, 2019. ISSN 1025-5834. URL https:/

/journalofinequalitiesandapplications.springeropen.com/
articles/10.1186/s13660-019-2107-6. Paper no. 159.

Gautschi:2003:EEI

[GHT03] W. Gautschi, F. E. Harris, and N. M. Temme. Expansions of the exponential integral in incomplete gamma functions. Applied Mathematics Letters, 16(7):1095–1099, October 2003. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Gautschi:1988:LBC

[GI88] Walter Gautschi and Gabriele Inglese. Lower bounds for the condition number of Vandermonde matrices. Numerische Mathematik, 52(3):241–250, March 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1970:RCC

[GK70a] Walter Gautschi and Bruce J. Klein. Recursive computation of certain derivatives — a study of error propagation. Communications of the ACM, 13(1):7–9, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1970:RAD

[GK70b] Walter Gautschi and Bruce J. Klein. Remark on Algorithm 282, Derivatives of e^{x}/x, cos(x)/x, and sin(x)/x. Communications of the ACM, 13(1):53–54, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gau66a].

Gautschi:1987:NES

[GKM87] Walter Gautschi, M. A. Kovačević, and Gradimir V. Milovanović. The numerical evaluation of singular integrals with coth-kernel. BIT (Nordisk tidskrift for informationsbehandling), 27(3):389–402, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=27&
issue=3&spage=389.

Gautschi:1990:RTA

[GL90] Walter Gautschi and Shikang Li. The remainder term for analytic functions of Gauss–Radau and Gauss–Lobatto quadrature rules with multiple end points. Journal of Computational and Applied Mathematics, 33(3):315–329, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037704270580007X.

Gautschi:1991:GRG

[GL91] Walter Gautschi and Shikang Li. Gauss–Radau and Gauss– Lobatto quadratures with double end points. Journal of Computational and Applied Mathematics, 34(3):343–360, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037704279190094Z.

Gautschi:1996:QCE

[GL96] Walter Gautschi and Shikang Li. On quadrature convergence of extended Lagrange interpolation. Mathematics of Computation, 65(215):1249–1256, July 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.org/

jourcgi/jour-pbprocess?fn=110&arg1=S0025-5718-96-00731-
4&u=/mcom/1996-65-215/.

Gautschi:2007:CIJ

[GL07] Walter Gautschi and Paul Leopardi. Conjectured inequalities for Jacobi polynomials and their largest zeros. Numerical Algorithms, 45(1–4):217–230, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.springer.

com/article/10.1007/s11075-007-9067-5.

Giordano:1998:UTG

[GLP98] C. Giordano, A. Laforgia, and J. Pecarić. Unified treatment of Gautschi–Kershaw type inequalities for the gamma function. Journal of Computational and Applied Mathematics, 99(1–2):167–175, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037704279800154X.

Gautschi:1977:OCQ

[GM77] Walter Gautschi and Giovanni Monegato. On optimal Chebyshevtype quadratures. Numerische Mathematik, 28(1):59–67, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1985:GQI

[GM85a] Walter Gautschi and Gradimir V. Milovanovi. Gaussian quadrature involving Einstein and Fermi functions with an application to summation of series. Mathematics of Computation, 44(169):177–190, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://www.jstor.org/stable/2007801.

Gautschi:1985:SGQ

[GM85b] Walter Gautschi and Gradimir V. Milovanovi. Supplement to Gaussian quadrature involving Einstein and Fermi functions with an application to summation of series. Mathematics of Computation, 44(169):S1–S11, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.jstor.org/

stable/2007816.

Gautschi:1986:SAS

[GM86] Walter Gautschi and Gradimir V. Milovanović. Spline approximations to spherically symmetric distributions. Numerische Mathematik, 49(2/3):111–121, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1997:OCG

[GM97] Walter Gautschi and Gradimir V. Milovanovic. S-orthogonality and construction of Gauss–Turán-type quadrature formulae. Journal of Computational and Applied Mathematics, 86(1):205–218, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042797001568.

Gautschi:2018:BTP

[GM18] Walter Gautschi and Gradimir V. Milovanović. Binet-type polynomials and their zeros. Electronic Transactions on Numerical Analysis, 50:52–70, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://etna.mcs.

kent.edu/vol.50.2018/pp52-70.dir/pp52-70.pdf;   http://
etna.mcs.kent.edu/volumes/2011-2020/vol50/abstract.php?
vol=50&pages=52-70.

Gautschi:2021:OPR

[GM21] Walter Gautschi and Gradimir V. Milovanović. Orthogonal polynomials relative to a generalized Marchenko–Pastur probability measure. Numerical Algorithms, 88(3):1233–1249, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01073-1.

Gautschi:2022:OPR

[GM22] Walter Gautschi and Gradimir V. Milovanović. Orthogonal polynomials relative to weight functions of Prudnikov type. Numerical Algorithms, 90(1):263–270, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01187-6.

Gautschi:2024:RID

[GM24] Walter Gautschi and Gradimir V. Milovanović. A Ramanujan integral and its derivatives: computation and analysis. Mathematics of Computation, 93(347):1297–1308, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.org/journals/mcom/2024-93-
347/S0025-5718-2023-03892-3.

Gautschi:2001:P

[GMR01] Walter Gautschi, Francisco Marcellán, and Lothar Reichel. Preface. Journal of Computational and Applied Mathematics, 127(1– 2):ix–xi, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0377042700004908.

Gautschi:2011:ACH

[GMR11] Walter Gautschi, Giuseppe Mastroianni, and Themistocles M. Rassias, editors. Approximation and Computation: In Honor of Gradimir V. Milovanović, volume 42 of Springer Optimization and Its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-6593-9 (paperback), 1-4419-6594-7 (e-book), 1-4419-6595-5, 1-4614-2703-7. xviii + 482 pp. LCCN QA39.2 .A67 2011; QA221 .A6345 2011.

Gautschi:1988:ASG

[GN88] Walter Gautschi and Sotirios E. Notaris. An algebraic study of Gauss–Kronrod quadrature formulae for Jacobi weight functions. Mathematics of Computation, 51(183):231–248, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1989:GKQ

[GN89] Walter Gautschi and Sotorios E. Notaris. Gauss–Kronrod quadrature formulae for weight functions of Bernstein–Szegő type. Journal of Computational and Applied Mathematics, 25(2):199–224, February 1989. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0377042789900472. See erratum [Ano89].

Gautschi:1996:SPR

[GN96] Walter Gautschi and Sotirios E. Notaris. Stieltjes polynomials and related quadrature formulae for a class of weight functions. Mathematics of Computation, 65(215):1257–1268, July 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00732-6&u=/mcom/1996-65-215/.

Golub:1984:SNA

[Gol84] Gene H. Golub, editor. Studies in Numerical Analysis, volume 24 of Studies in mathematics. Mathematical Association of America, Washington, DC, USA, 1984. ISBN 0-88385-126-1 (v. 1), 0-88385-100-8 (set). x + 415 pp. LCCN QA297 .S83 1984.

Gautschi:1988:FGK

[GR88] Walter Gautschi and Theodore J. Rivlin. A family of Gauss– Kronrod quadrature formulae. Mathematics of Computation, 51 (184):749–754, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grimm:2005:NGT

[Gri05a] Volker Grimm. A note on the Gautschi-type method for oscillatory second-order differential equations. Numerische Mathematik, 102(1):61–66, November 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=102&issue=1&spage=61.

Grimm:2005:EBG

[Gri05b] Volker Grimm. On error bounds for the Gautschi-type exponential integrator applied to oscillatory second-order differential equations. Numerische Mathematik, 100(1):71–89, March 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1978:CMB

[GS78] Walter Gautschi and Josef Slavik. On the computation of modified Bessel function ratios. Mathematics of Computation, 32(143):865– 875, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1990:NCI

[GTV90] Walter Gautschi, E. Tychopoulos, and R. S. Varga. A note on the contour integral representation of the remainder term for a Gauss– Chebyshev quadrature rule. SIAM Journal on Numerical Analysis, 27(1):219–224, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gautschi:1983:EBG

[GV83] Walter Gautschi and Richard S. Varga. Error bounds for Gaussian quadrature of analytic functions. SIAM Journal on Numerical Analysis, 20(6):1170–1186, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gautschi:1984:MYL

[GW84] Walter Gautschi and Jet Wimp. In memoriam: Yudell L. Luke, June 26, 1918–May 6, 1983. Mathematics of Computation, 43 (168):349–352, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1987:CHT

[GW87] Walter Gautschi and Jet Wimp. Computing the Hilbert transform of a Jacobi weight function. BIT (Nordisk tidskrift for informationsbehandling), 27(2):203–215, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=27&issue=2&spage=203.

Gautschi:2001:CHT

[GW01] Walter Gautschi and Jörg Waldvogel. Computing the Hilbert transform of the generalized Laguerre and Hermite weight functions. BIT Numerical Mathematics, 41(3):490–503, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125

(electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=41&issue=3&spage=490.

Gautschi:1974:CTQ

[GY74] Walter Gautschi and Hiroki Yanagiwara. On Chebyshev-type quadratures. Mathematics of Computation, 28(125):125–134, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1995:COP

[GZ95] Walter Gautschi and Minda Zhang. Computing orthogonal polynomials in Sobolev spaces. Numerische Mathematik, 71(2): 159–183, August 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.de/

link/service/journals/00211/bibs/5071002/50710159.htm;
 http://science.springer.de/nmee/bibs/5071002/50710159.
htm.

Higham:1994:SCPb

[Hig94] Nicholas J. Higham. A survey of componentwise perturbation theory in numerical linear algebra. In Walter Gautschi, editor, Mathematics of Computation 1943–1993: A Half Century of Computational Mathematics, volume 48 of Proceedings of Symposia in Applied Mathematics, pages 49–77. American Mathematical Society, Providence, RI, USA, 1994.

Hochbruck:1999:GTM

[HL99] Marlis Hochbruck and Christian Lubich. A Gautschi-type method for oscillatory second-order differential equations. Numerische Mathematik, 83(3):403–426, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer-ny.com/link/service/journals/00211/bibs/
9083003/90830403.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/9083003/90830403.pdf.

I:1991:BRH

[I.91] E. I. Book review: Heinz Rutishauser, Lectures on Numerical Mathematics (translated from the German by Walter Gautschi), Birkhauser, Boston, 1990, xv + 546 pp., 23.5 cm. Price $49. Mathematics of Computation, 57(196):869–870, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.jstor.org/stable/2938724.

Jansen:1977:RLF

[Jan77] J. K. M. Jansen. Remark on “Algorithm 259: Legendre functions for arguments larger than one”. ACM Transactions on Mathematical Software, 3(2):204–205, June 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Gau65a].

Joulak:2009:GCG

[JB09] Hédi Joulak and Bernhard Beckermann. On Gautschi’s conjecture for generalized Gauss–Radau and Gauss–Lobatto formulae. Journal of Computational and Applied Mathematics, 233(3):768–774, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042709001241.

Jeltsch:2008:ICI

[JW08] Rolf Jeltsch and Gerhard Wanner, editors. 6th International congress on industrial and applied mathematics, ICIAM 07, Zürich, Switzerland, 16-20.07.2007. European Mathematical Society Publishing House, Zürich, Switzerland, 2008. ISBN 3-03719-056-6, 3-03719-556-8 (e-book). LCCN QA1 .I73 2007.

Kahan:196x:ACM

[Kah6x] W. M. Kahan. Note on bounds for generating Bessel functions by recurrence. 196x.

Kolm:1971:MAL

[KD71] Å. Kolm and T. Dahlstrand. Remark on “Algorithm 333: Minit Algorithm For Linear Programming ([H])”. Communications of the ACM, 14(1):50, 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gau64a].

Kershaw:1983:SEW

[Ker83] D. Kershaw. Some extensions of W. Gautschi’s inequalities for the gamma function. Mathematics of Computation, 41(164):607–611, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1972:CAC

[Köl72] K. S. Kölbig. Certification of “Algorithm 363: Complex error function”. Communications of the ACM, 15(6):465–466, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gau69b].

Koumandos:2007:CIG

[Kou07] Stamatis Koumandos. On a conjectured inequality of Gautschi and Leopardi for Jacobi polynomials. Numerical Algorithms, 44 (3):249–253, March 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=44&
issue=3&spage=249.

Kunisch:2015:GTS

[KR15] Karl Kunisch and Stefan H. Reiterer. A Gautschi time-stepping approach to optimal control of the wave equation. Applied Numerical Mathematics: Transactions of IMACS, 90(??):55–76, April 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0168927414001937.

Lenstra:1997:BRM

[Len97] H. W. Lenstra, Jr. Book reviews: Mathematics of Computation 1943-1993: A half-century of computational mathematics, by Walter Gautschi (Editor). Mathematics of Computation, 66(219):??, July 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lide:2001:CEM

[Lid01] D. R. Lide, editor. A Century of Excellence in Measurements, Standards, and Technology: A Chronicle of Selected NBS/NIST Publications, 1901–2000, volume 958. National Technical Information Service, Washington, DC, USA, 2001. ix + 386

pp. URL https://nvlpubs.nist.gov/nistpubs/sp958-lide/

cntsp958old.htm. NIST Special Publication.

Lun:2014:IZJ

[LR14] Yen Chi Lun and Fernando Rodrigo Rafaeli. Inequalities for zeros of Jacobi polynomials via Sturm’s theorem: Gautschi’s conjectures. Numerical Algorithms, 67(3):549–563, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11075-013-9807-7.

Pomerance:1994:NFS

[Pom94] C. Pomerance. The number field sieve. In W. Gautschi, editor, Mathematics of Computation 1943–1993: A Half-Century of Computational Mathematics. Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of symposia in applied mathematics, pages 465–480. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). LCCN QA1 .A56 v.48 1994.

Pozza:2018:LAC

[PPS18] Stefano Pozza, Miroslav S. Pranić, and Zdeněk Strakoš. The Lanczos algorithm and complex Gauss quadrature. Electronic Transactions on Numerical Analysis, 50:1–19, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://etna.

mcs.kent.edu/vol.50.2018/pp1-19.dir/pp1-19.pdf;   http:
//etna.mcs.kent.edu/volumes/2011-2020/vol50/abstract.php?
vol=50&pages=1-19.

Qi:2008:WGI

[QG08] Feng Qi and Bai-Ni Guo. Wendel’s and Gautschi’s inequalities: Refinements, extensions, and a class of logarithmically completely monotonic functions. Applied Mathematics and Computation, 205(1):281–290, November 1, 2008. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300308005377.

Qi:2009:CLC

[Qi09] Feng Qi. A class of logarithmically completely monotonic functions and application to the best bounds in the second Gautschi– Kershaw’s inequality. Journal of Computational and Applied Mathematics, 224(2):538–543, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042708002604.

Rutishauser:1990:LNM

[RG90] Heinz Rutishauser and Martin Gutknecht, editors. Lectures on numerical mathematics. Birkhäuser Boston Inc., Cambridge, MA, USA, 1990. ISBN 0-8176-3491-6. xv + 546 pp. LCCN QA297 .R8713 1990. Edited by Martin Gutknecht with the assistance of Peter Henrici, Peter Läuchli and Hans-Rudolf Schwarz, With a foreword by Gutknecht and a preface by Henrici, Läuchli and Schwarz, Translated from the German and with a preface by Walter Gautschi.

Rice:1994:FBW

[Ric94] John R. Rice. Foreword [on the 65th birthday of Walter Gautschi]. In Zahar [Zah94b], pages ix–x. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Rojas:1994:WIC

[Roj94] Raúl Rojas. Who invented the computer? The debate from the viewpoint of computer architecture. In Walter Gautschi, editor, Mathematics of Computation 1943–1993: A Half Century of Computational Mathematics, volume 48 of Proceedings of Symposia in Applied Mathematics, pages 361–365. American Mathematical Society, Providence, RI, USA, 1994.

Stoer:1993:INA

[SB93] Josef Stoer and Roland Bulirsch. Introduction to numerical analysis, volume 12 of Texts in Applied Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1993. ISBN 0-387-97878-X (New York), 3-540-97878-X (Berlin). xiii + 660 pp. LCCN QA297 .S8213 1993. Translated from the German by R. Bartels, W. Gautschi and C. Witzgall.

Schneider:1981:HAM

[SDGT81] Hans Schneider, Chandler Davis, Walter Gautschi, and Olga Taussky Todd. Honoring Alexander M. Ostrowski. Linear Algebra and its Applications, 38(??):3, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900021.

Segura:2006:ROP

[Seg06] Javier Segura. Review of Orthogonal Polynomials: Computation and Approximation by Walter Gautschi. SIAM Review, 48(2): 431–433, June 2006. CODEN SIREAD. ISSN 0036-1445 (print),

1095-7200 (electronic). URL https://www.jstor.org/stable/
20453824.

Skovgaard:1975:RBF

[Sko75] Ove Skovgaard. Remark on “Algorithm 236: Bessel Functions of the First Kind [S17]”. ACM Transactions on Mathematical Software, 1 (3):282–284, September 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Gau64b].

Stetter:1999:BRN

[Ste99] Hans J. Stetter. Book reviews: Numerical analysis, an introduction, by Walter Gautschi. Mathematics of Computation, 68(226):887, April 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.org/

jourcgi/jour-pbprocess?fn=110&arg1=S0025-5718-99-01060-
1&u=/mcom/1999-68-226/.

Shi:2019:EGT

[SW19] Wei Shi and Xinyuan Wu. Explicit Gautschi-type integrators for nonlinear multi-frequency oscillatory second-order initial value problems. Numerical Algorithms, 81(4):1275–1294, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Townsend:2016:ROP

[Tow16] Alex Townsend. A review of Orthogonal Polynomials in MATLAB: Exercises and Solutions by Gautschi. Web document., Novem-

ber 5, 2016. URL http://pi.math.cornell.edu/~ajt/papers/
Gautschi.pdf.

Weideman:1995:ECE

[Wei95] J. A. C. Weideman. Erratum: “Computation of the complex error function” [SIAM J. Numer. Anal. 7 (1970), no. 1, 187–198]. SIAM Journal on Numerical Analysis, 32(1):330–331, February 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Gau70a].

Wong:1990:ACA

[Won90] R. (Roderick) Wong, editor. Asymptotic and Computational Analysis: Conference in Honor of Frank W. J. Olver’s 65th Birthday, volume 124 of Lecture Notes in Pure and Applied Mathematics. Marcel Dekker, New York, NY, USA, 1990. ISBN 0-8247-8347-6.

LCCN QA299.6 .A88 1990. URL http://www.loc.gov/catdir/
enhancements/fy0647/90002810-d.html.

Xu:2016:EEE

[XDY16] Zhiguo Xu, Xuanchun Dong, and Yongjun Yuan. Error estimates in the energy space for a Gautschi-type integrator spectral discretization for the coupled nonlinear Klein–Gordon equations. Journal of Computational and Applied Mathematics, 292(??):402–416, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042715003799.

Zahar:1994:IBW

[Zah94a] R. V. M. Zahar. Introduction [on the 65th birthday of Walter Gautschi]. In Zahar [Zah94b], pages xi–xvi. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Zahar:1994:ACF

[Zah94b] R. V. M. (Ramsay Vincent Michael) Zahar, editor. Approximation and computation: a Festschrift in honor of Walter Gautschi: proceedings of the Purdue conference, December 2–5, 1993, volume 119 of International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Zahar:1994:PWG

[Zah94c] R. V. M. (Ramsay Vincent Michael) Zahar. Publications of Walter Gautschi. In Approximation and computation: a Festschrift in honor of Walter Gautschi: proceedings of the Purdue conference, December 2–5, 1993 [Zah94b], pages xxxvi–xlvi. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.